### v1

# Towards Effectively Detecting and Explaining Vulnerabilities Using Large Language Models

Abstract.
Software vulnerabilities pose significant risks to the security and integrity of software systems. Prior studies have proposed a series of approaches to vulnerability detection using deep learning or pre-trained models. However, there is still a lack of vulnerability’s detailed explanation for understanding apart from detecting its occurrence. Recently, large language models (LLMs) have shown a remarkable capability in the comprehension of complicated context and content generation, which bring opportunities for the detection and explanation of vulnerabilities of LLMs. In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in detecting and explaining vulnerabilities and propose LLMVulExp, a framework that utilizes LLMs for vulnerability detection and explanation. Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities. We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., over 90% F1 score on SeVC dataset) and explanation. We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs concentrating on vulnerability-prone code and achieve promising results.
Co-first authors, equally contributed.
Corresponding author.
copyright: noneconference: ; June 2024; Online
1. Introduction
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions (Zhou et al., 2019; Li et al., 2021a). These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
Current vulnerability detection techniques mainly include pattern based and deep learning based approaches. Pattern based approaches (Yamaguchi, 2017; rou, [n. d.]) generally rely on manually defined rules to detect vulnerabilities. Deep learning based approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) train the models using existing vulnerability data and various code representation techniques. Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities. This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage. It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations. Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue. LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations. However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation. They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection. This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
Figure 1. Example of vulnerability detection result with and without explanation.
The challenges of this fine-tuning process are twofold: obtaining high-quality training data (Zhou et al., 2019; Chen et al., 2023; Zhou et al., 2024; Zhao et al., 2023; Wang et al., 2024) and effectively adapting the models (Xu et al., 2023; Hu et al., 2021; Thapa et al., 2022). Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial. These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws. Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation. Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
In this study, we explore enhancing LLMs to detect and explain software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using proprietary LLM models. The annotated explanatory information includes the location of the vulnerability and a detailed explanation. Following this, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance (Fu et al., 2023a). Additionally, we investigate the correlation between vulnerability detection and explanation, as well as utilizing LLMs to follow a Chain-of-Thought (CoT) (Wei et al., 2022) strategy and perform a concentrated vulnerability detection on the key code that is prone to software vulnerabilities. Furthermore, we address the evaluation of generated vulnerability explanations (Cheshkov et al., 2023) by proposing new evaluation metrics and an automated review method using LLMs.
Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC (Li et al., 2021c) and DiverseVul (Chen et al., 2023)), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation. We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
Overall, the contributions of this paper are threefold:
  • We present a comprehensive and effective workflow for training, inferring, and evaluating LLMs based on open-source vulnerability data.
  • We pioneer the exploration of the effectiveness of vulnerability explanation models and propose novel evaluation methods and dimensions for these explanations.
  • We analyze the interrelationship between explanatory and detection capabilities, offering valuable insights for the development of explainable vulnerability detection methodologies.
Paper Organization. Section 2 summarizes the related work. Section 3 presents the methodology of our study. Section 4 discusses the results to our research questions. Section 5 discusses the implications of our study. Section 6 discusses the threats to validity. Section 7 concludes the paper.
2. Related Work
In this section, we summarize the related work on vulnerability detection and explanation, respectively.
2.1. Vulnerability Detection
Prior studies proposed a series of deep learning approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) to detect the vulnerabilities. These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
Apart from deep learning approaches, recent advancements in large language models (LLMs) have significantly influenced the field of software vulnerability detection. Advanced methodologies such as zero-shot prompting (Zhou et al., 2024; Fu et al., 2023b), in-context learning (Dong et al., 2022; Brown et al., 2020a), and fine-tuning (Hu et al., 2021; Xu et al., 2023; Chen et al., 2023) have been employed in vulnerability detection. Cheshkov et al. (Cheshkov et al., 2023) evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings. Gao et al. (Gao et al., 2023) proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods. Nong et al. (Nong et al., 2024) focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting. By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy. Sun et al. (Sun et al., 2024) further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities. They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance. Yusuf et al. (Yusuf and Jiang, 2024) discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages. Steenhoek et al. (Steenhoek et al., 2024) surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
These studies underscore the challenges and opportunities of applying LLMs to vulnerability detection. Unlike prior methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
2.2. Vulnerability Explanation
Although significant research progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these techniques for vulnerability explanation remains a challenging issue. Currently, only a limited number of studies focus on the explanatory capabilities of deep learning based models regarding vulnerabilities.
The VulDeeLocator  (Li et al., 2021b) enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer. In a similar vein, both IVDetect (Li et al., 2021a) and LineVul (Fu and Tantithamthavorn, 2022) develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements. VELVET (Ding et al., 2022) prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks. LineVD (Hin et al., 2022) addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning. VulTeller (Zhang et al., 2023) emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities. VulExplainer  (Cheng et al., 2024) is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks. Coca (Cao et al., 2024) is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
The semantic understanding of neural network offers the potential for vulnerability explanation models tailored to software development. However, accurate detection and effective explanation of vulnerabilities using LLMs remain substantial challenges for general-purpose code models. Our work explores the feasibility of fine-tuning specialized LLMs for vulnerability detection and explanation, addressing a crucial gap in this research domain.
3. Methodology
Figure 2. Overview of our study.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to identify and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for vulnerability detection and explanation. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important. In this paper, we conduct the study on two datasets: (1) SeVC (Li et al., 2021c), which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul (Chen et al., 2023), which covers 295 real open-source projects and 150 CWE types.
Table 1. Statistics of the studied datasets.
Dataset Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
SeVC 56,395 40,491 4 Single/Multi-Type
DiverseVul 18,945 9,161 10 Multi-Type(CWE)
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 Common Weakness Enumerations (CWEs). DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise. Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types. We select the top ten CWE types to construct our fine-tuning dataset, enabling us to conduct multi-label classification detection
We deduplicate the vulnerability samples using a hash method. Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset. The dataset details are shown in Table 1. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets primarily include information within the scope of source code, vulnerability labels, CWE types, and commit messages. However, there is still a lack of detailed explanations of the vulnerability logic within the source code, which poses challenges for vulnerability detection techniques to provide a corresponding explanation regarding the detection results. Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, involving high labor and time costs.
To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using open-source LLMs. This method leverages the contextual learning and instruction-following capabilities of LLMs, using prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation. Prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code localization, and specific explanation. Combining instruction-based prompt templates and effective annotated examples to stimulate the model’s contextual learning capabilities ensures the effectiveness of vulnerability explanation annotation.
In this paper, we use GPT-3.5 (Brown et al., 2020b) by invoking the API provided by OpenAI (ope, 2024) to implement the annotation process. To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills the current gap in vulnerability explanation data. The specific annotation process can be found in Section 4.1.
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data provides a large-scale dataset for vulnerability detection and explanation, addressing the data bottleneck issues in the fine-tuning process. To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune general-purpose LLMs to enable them to detect and explain specific types of vulnerabilities in real code. We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized output. To reduce the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021), significantly reducing time and space costs.
❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations makes effective evaluation of LLM-generated vulnerability explanations an important issue. Similar to the challenges faced in annotation, manual evaluation also requires significant human and time resources. Furthermore, the effectiveness of vulnerability explanations involves multiple dimensions to provide useful assistance to developers in real development scenarios. To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and constructiveness, combined with prompt engineering to develop an automated LLM evaluation method. Through expert manual verification, we validate the effectiveness of the specialized vulnerability explanation model’s generated quality and the feasibility of the LLM-based automated evaluation scheme.
3.1. Vulnerability Interpretation Enhancement Prompting
Despite the strong code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning, especially in tasks that require a deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities. These challenges manifest as insufficient detection accuracy and vague vulnerability analysis. To guide LLMs towards better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques and the contextual learning capabilities of LLMs. By combining these with prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
Figure 3 shows the design of our prompt templates. The prompt templates consist of four main components: task description, specific instructions, generation examples, and sample inputs. (1) Task description: provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format, helping the LLM understand task requirements and grasp the background knowledge of the vulnerabilities. (2) Specific instructions: include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output. These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis. (3) Generation examples: provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals. (4) Sample input: is used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
Figure 3. Prompting template of Automated Vulnerability Interpretation Labelling.
Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, leading to well-performing fine-tuned specialized vulnerability detection and explanation models. This confirms the effectiveness of our prompt design approach.
3.2. Key Code Extraction Based on Chain-of-Thought (CoT)
One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets. Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs, but these methods are difficult to apply in diverse vulnerability detection scenarios. To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction. This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code. The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities in the form of a thinking chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
Figure 4 illustrates the prompt template for key statement extraction designed for LLMs. The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on. After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT. This guides the model in reference to the key statements step by step to complete the vulnerability detection, localization, and explanation.
Figure 4. Prompting template of Key Code Extraction.
4. Results
In this section, we discuss the results by proposing and answering the following research questions:
  • RQ1: How effective are LLMs in detecting software vulnerabilities?
  • RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
  • RQ3: How do explanations affect the results of vulnerability detection?
  • RQ4: How does the key code extraction impact detection performance?
4.1. RQ1: How effective are LLMs in detecting software vulnerabilities?
In this RQ, we discuss the detection performance of fine-tuned, vulnerability-specialized LLMs in various scenarios.
4.1.1. Experimental Setup.
Dataset. As discussed in Section 3, we select the SeVC (Li et al., 2021c) dataset and the DiverseVul dataset (Chen et al., 2023). assessments involving ten CWE types of vulnerabilities.
Backbone LLMs. We use the Codellama (Roziere et al., 2023) and Llama3 as our base models for fine-tuning. CodeLlama is initialized with the weights of Llama2 (Touvron et al., 2023) and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. Considering the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct (cod, 2023) as our primary model and used the 7B version and the newly released Llama3-8B-Instruct for comparison (refer to the discussion section).
Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability models, we design four sub-research questions:
  • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
  • (RQ1.2) What impact does the inclusion of other types of vulnerability data in training have on the detection performance of the target type of vulnerability?
  • (RQ1.3) How effective is a unified model in detecting multiple types of vulnerabilities?
  • (RQ1.4) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenarios in RQ1.1 and RQ1.2, we directly use these three metrics. For the multi-class classification in RQ1.3, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.4, we use Micro-F1 and Macro-F1. To test the detection accuracy, we selected CodeT5 (Wang et al., 2021) and CodeBERT (Feng et al., 2020) as baselines and fine-tuned them for classification tasks by adding a linear classification layer.
Implementation Details. We implement our approach using the Transformers (Thapa et al., 2022) and PEFT (Mangrulkar et al., 2022) libraries with the PyTorch platform. All experiments are conducted on NVIDIA A100-SXM4-80GB GPU platforms, with the token length limit set to 2048. We use the AdamW optimizer and train the models for 3 epochs. For Low-rank Adaptation (LoRA) configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
4.1.2. RQ1.1: Detection with specified vulnerability type using a dedicated model.
To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type. Specifically, we indicate the type of vulnerability we are concerned with in the training and inference prompts to help it narrow down the scope of detecting vulnerabilities. During training, we use the explanations with vulnerability location information and specific details annotated by GPT as the target output for the model to perform detection and explanation tasks. The inputs for the model during training and inference are code snippets of the given examples. During inference, we develop a rule to analyze the generated result to determine whether it indicates a vulnerability or not. Specifically, we consider a result as non-vulnerable if it contains the sentences following a fixed pattern (e.g., “There is no security issues” and “detected in the given code snippet.”). Otherwise, we consider the result indicates a vulnerability. We randomly selected 100 vulnerable results and 100 non-vulnerable results to examine the correctness of such rule-based analysis. We find that all of the vulnerable and non-vulnerable results can be correctly classified by this rule.
Table 2. Performances of Single-Type Detection on SeVC(RQ1.1).
Precision Recall F1
Type Ours CodeT5 CodeBert Ours CodeT5 CodeBert Ours CodeT5 CodeBert
API 91.6% 92.2% 93.2% 94.5% 88.1% 87.1% 93.0% 90.1% 90.0%
Arithmetic 90.3% 88.2% 90.9% 93.6% 94.7% 98.0% 91.9% 91.3% 91.9%
Pointer 93.7% 93.5% 95.8% 91.2% 95.0% 93.4% 92.4% 94.3% 94.6%
Array 95.3% 92.9% 95.1% 89.7% 94.1% 92.2% 92.4% 93.5% 93.6%
Average 92.7% 91.7% 93.7% 92.3% 93.0% 92.7% 92.4% 92.3% 92.2%
Experimental Results. The experimental results are shown in Table 2. The fine-tuned models on four types of vulnerabilities have exceeded 90% performances on three metrics, which indicates that the specialized models for a single type of vulnerability can achieve good results. On one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type. On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs. If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
4.1.3. RQ1.2: Detection with specified vulnerability type using an all-in-one model.
To explore the impact of training with data from other types of vulnerabilities on the detection effect of the target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning on four types of vulnerabilities and performing individual detection for each type. Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection, and during inference, we also provide the vulnerability type information for each example. The difference between RQ1.1 and RQ1.2 is that, we fine-tune a dedicated model for each type of vulnerability in RQ1.1. In RQ1.2, we fine-tune an all-in-one model and use the prompts to distinguish the type of vulnerability in interest.
Table 3. Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
Accuracy Precision Recall F1
Type Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one
API 93.0% 97.1% 91.6% 97.4% 94.5% 96.8% 93.0% 97.1%
Arithmetic 91.7% 96.6% 90.3% 97.6% 93.6% 95.6% 91.9% 96.6%
Pointer 92.7% 97.1% 93.7% 97.6% 91.2% 96.6% 92.4% 97.1%
Array 92.7% 96.6% 95.3% 97.0% 89.7% 96.3% 92.4% 96.6%
Average 92.5% 96.9% 92.7% 97.4% 92.2% 96.3% 92.4% 96.9%
Experimental Results. We present the results of the fine-tuned all-in-one model in Table 3. In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities. This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities. This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs. Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
4.1.4. RQ1.3: Detection with an identification of the vulnerability type.
In real software development environments, the vulnerability risks faced are often diverse. It may not always be possible to preemptively sense the potential types of vulnerabilities in the samples to be detected. This requires the large model to correctly identify the type of vulnerabilities present from multiple types and further provide specific explanations. To test the accuracy of a unified model in identifying and classifying code with multiple types of vulnerabilities, we conduct a multi-class vulnerability detection fine-tuning on SeVC. Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including none-vulnerable and the four types of vulnerabilities in SeVC. In the training and inference prompts, we no longer specify the type of vulnerability, but instead turn to a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
Table 4. Performances of Multi-Type Detection on SeVC(RQ1.3).
Precision Recall F1 Score
Type Ours CodeT5 CodeBert Ours CodeT5 CodeBert Ours CodeT5 CodeBert
Non-vul 95.4% 94.7% 94.6% 98.0% 94.6% 96.7% 96.7% 94.7% 95.7%
Array 61.9% 56.0% 58.4% 55.2% 65.6% 62.8% 58.3% 60.4% 60.5%
Pointer 72.9% 72.2% 74.6% 70.9% 66.8% 70.9% 71.9% 69.4% 70.0%
API 48.8% 45.4% 44.5% 46.0% 40.8% 43.5% 47.4% 43.0% 44.0%
Arithmetic 70.7% 65.3% 68.3% 91.6% 88.6% 87.3% 79.8% 75.2% 76.7%
Weighted 79.9% 78.2% 78.9% 80.5% 78.1% 79.1% 80.1% 78.0% 79.1%
Macro 70.0% 66.7% 68.1% 72.4% 71.3% 71.3% 70.8% 68.5% 69.4%
Experimental Results. We present the metrics for each type and the confusion matrix for the multi-class vulnerability detection task in Table 4. Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code. (2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario. (3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types. The other three types exhibit a significant degree of confusion and are difficult to identify. This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient. The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
4.1.5. RQ1.4: Identification of the vulnerability type on the datasets with more vulnerability types.
In real development environments, the security risks of vulnerabilities come from a greater variety of project types and more complex code structures, often involving a wider range of vulnerability models. To more closely resemble real development conditions, we conduct a multi-label classification task of the top 10 types of CWE on the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels). Specifically, we modify the task description to be based on CWE-type vulnerability detection and explanation tasks, and add the ‘[CWE]’ tag for generating a list of CWEs related to the target code. At the same time, we incorporate the CWE descriptions as part of the output explanation information to help the model further understand the meaning of each CWE type of vulnerability and to use the CWE descriptions for vulnerability analysis and identification.
Table 5. Performances of Multi-Label CWE Detection on DiverseVul(RQ1.4).
Metric CWE ID Overall
787 125 119 20 703 416 476 190 200 399 Micro Macro
Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 73.3%
Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 73.4%
F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 73.2%
Experimental Results. The experimental results are presented in Table 5. From the results, we find that after fine-tuning, the specialized vulnerability models have achieved good detection accuracy for the 10 types of CWE, demonstrating an ability to grasp the pattern information for each CWE type. Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories. Each vulnerability type has a more fine-grained explicit definition, and it also avoids the potential category association issues that might exist in SeVC, thus resulting in better performance. This validates the feasibility of research into LLMs for vulnerability detection and explanation that are oriented toward real software development environments.
Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
4.2. RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities. Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanation. The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content. Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics, Accuracy, Clarity, and Constructiveness, as the evaluation criteria for vulnerability explanation. We use manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
4.2.1. Experimental Setup.
In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT. We randomly sample the data based on 95% confidence level and 5% confidence interval (Boslaugh and Watters, 2008) for the fine-tuned model’s correctly classified instances, obtaining 384 and 224 samples from SeVC and DiverseVul, respectively. Our evaluation includes two parts, (1) Manual review: two authors of this paper independently conduct manual checks on the samples, scoring them according to the criteria. In cases of disagreement, discussions are held until a consensus. The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement. (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion. We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction. This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.
Evaluation Metrics. The specific meanings and requirements of the three metrics are as follows:
  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring that the provided details are factually correct and relevant to the detected vulnerability.
  • Clarity: The information should be presented in a clear, understandable manner, and structured in a way that facilitates easy comprehension by software developers.
  • Constructiveness: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
4.2.2. Experimental Results.
Table 6. Distribution (%) of Vulnerability Explanation Review Results (RQ2).
Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
Manual
Accuracy 91.1 93.1 73.3 94.1
Clarity 81.4 81.7 94.1 98.6
Construct. 93.4 94.6 80.5 83.2
All-Pos. 76.0 76.0 59.7 80.1
LLM-Automation
Accuracy 90.4 97.6 96.4 96.8
Clarity 74.3 77.2 95.0 96.8
Construct. 83.5 88.6 67.9 71.9
All-Pos. 72.8 74.0 67.9 71.9
We present the results of our dual evaluation process in Table 6, the suffix Gen. refers to the explanation generated by LLMVulExp and Ann. refers to the explanation annotated by GPT. All-Pos. refers to the results that meet requirements of all of the three metrics. Combining the results from manual and LLM evaluations, we have made the following observations:
(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen. for both the manual review and automated review) in the explanation. They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively. The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
(3) Constructiveness of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann. and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data. It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review. For SeVC and DiverseVul, the results of All-Pos. comparing manual review and automated review are 76.0% vs. 74.0% and 80.1% and 71.9%, respectively. This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text. In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation. The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations. The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities. The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis. Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Constructiveness. We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
4.3. RQ3: How do explanations affect the results of vulnerability detection?
By considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1. This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs, compared to fine-tuning specialized detection models? In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection (i.e., comparing the performance of vulnerability detection between fine-tuned LLMs with and without the explanatory information).
Experimental Setup. We conduct our research in the form of ablation studies under the settings of multi-type vulnerability detection scenario in SeVC (RQ1.3) and multi-label CWE detection scenario in DiverseVul (RQ1.4). We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt. By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model. We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
Figure 5. Performance Comparison between Explainer and Detector.
Experimental Results. We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 5, the prefix of “Mi” and “Ma” refers to Micro and Macro. Based on the experimental results, incorporating the vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance. For example, the weightd precision on SeVC for Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% and 74.8% on DiverseVul. In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities. The model’s detection capability still remains with an additional focus on explanation capabilities. Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns. It suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance. Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
Summary of RQ3: Integrating vulnerability explanation into the fine-tuning process does not compromise detection capabilities and may even lead to improved performance for certain vulnerability scenarios.
4.4. RQ4: How does the key code extraction impact detection performance?
In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to the vulnerability and further examine the code with a specific focus. To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
Experimental Setup. Similar to the process of annotation discussed in Section 3, we use GPT-3.5 to extract the key code. The prompt templates are illustrated in Figure 4. To prevent label leakage, the code vulnerability type tags were not visible during the annotation process. For training data that does not output the original code statements, we will nullify its key code information. If it is test data, it will not be included in the evaluation. During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
Table 7. F1 Scores of Key Code Extraction on SeVC (RQ4)
Type Single-Type(RQ1.1) Multi-Type(RQ1.3)
W/O Key. With Key. W/O Key. With Key.
API 93.0% 98.7% 79.8% 87.8%
Arith. 91.9% 97.9% 71.9% 96.0%
Pointer 92.4% 99.1% 58.3% 73.5%
Array 92.4% 98.2% 47.4% 77.9%
Average 92.4% 98.5% 64.4% 83.8%
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table 7 and DiverseVul in Figure 6. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results. On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio. This makes it easier for the model to learn and detect patterns associated with vulnerabilities.
Figure 6. Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent for the CWE IDs.
Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., 64.4% -¿ 83.8% for Multi-Type detection on SeVC).
5. Discussion
In this section, we discuss the implications of our study.
Implication 1: Implications for Broader Vulnerability-Related Tasks. Our study reveals that integrating the task of vulnerability explanation with detection does not compromise the model’s performance. This finding suggests we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs. By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights. For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime. This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs (i.e., CodeLlama-7B-Instruct and Llama3-8B-Instruct) to conduct the experiments following the setting of RQ1.3. Figure 7 shows the results of Multi-Type vulnerability detection using LLMs with different sizes. We find that LLMVulExp can also achieve effective performance using the additional LLMs. Particularly, the performance between CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct. Overall, LLMVulExp presents its effectiveness using different LLMs with various sizes.
Figure 7. Performance of Multi-Type vulnerability detection using LLMs with different sizes.
Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks. The effectiveness of fine-tuning heavily depends on the quality of the annotations used for training. Our study highlights the critical role that high-quality annotations play in the fine-tuning process. It is essential to develop robust annotation frameworks that ensure consistency and accuracy. Furthermore, the annotation process itself could be enhanced through the use of semi-automated tools that provide initial labels, which are then reviewed and refined by human experts. This hybrid approach could balance the need for detailed annotations and the practical constraints of manual labeling.
6. Threats to Validity
Construct Validity. We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Constructiveness. These metrics characterize three aspects of effective vulnerability explanations but may not comprehensively cover all features of vulnerability explanations. More comprehensive and refined evaluation metrics for vulnerability explanations require further research. In this paper, we employed both manual review and automated review using LLMs. To mitigate the bias, two authors independently annotated the explanations and resolved any discrepancies until a consensus. The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conducted at least two or more trials to validate the stability of our conclusions for each experimental setup. The number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited. Additionally, due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version. Although these limitations may have resulted in alternative findings, they do not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models. We conducted experiments using two differently configured open-source project-based vulnerability datasets. While these datasets contain a lot of vulnerable code, they differ from real-world development environments. Therefore, our findings may not apply to other datasets or industrial settings.
7. Conclusion
In this paper, we propose LLMVulExp, a framework that aims that detecting and explaining vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security. Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It suggests that a multi-task learning approach could be beneficial, addressing the data volume bottleneck is crucial for practical applications, and the quality of annotations is paramount for the success of fine-tuning frameworks. By considering these insights, future work can aim to develop more capable and efficient models that contribute to the field of software security.

### v2

@@
-Towards Effectively Detecting and Explaining Vulnerabilities Using Large Language Models
-Abstract.
+mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt
+Abstract
 Software vulnerabilities pose significant risks to the security and integrity of software systems.
-Prior studies have proposed a series of approaches to vulnerability detection using deep learning or pre-trained models.
-However, there is still a lack of vulnerability’s detailed explanation for understanding apart from detecting its occurrence.
-Recently, large language models (LLMs) have shown a remarkable capability in the comprehension of complicated context and content generation, which bring opportunities for the detection and explanation of vulnerabilities of LLMs.
-In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in detecting and explaining vulnerabilities and propose LLMVulExp, a framework that utilizes LLMs for vulnerability detection and explanation.
+Prior studies have proposed various approaches to vulnerability detection using deep learning or pre-trained models.
+However, there is still a lack of detailed explanations for understanding vulnerabilities beyond merely detecting their occurrence, which fails to truly help software developers understand and remediate the issues.
+Recently, large language models (LLMs) have demonstrated remarkable capabilities in comprehending complex contexts and generating content, presenting new opportunities for both detecting and explaining software vulnerabilities.
+In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in both detecting and explaining vulnerabilities, and we propose LLMVulExp, a framework that utilizes LLMs for these tasks.
 Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities.
-We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., over 90% F1 score on SeVC dataset) and explanation.
-We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs concentrating on vulnerability-prone code and achieve promising results.
-Co-first authors, equally contributed.
-Corresponding author.
-copyright: noneconference: ; June 2024; Online
-1. Introduction
-A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions (Zhou et al., 2019; Li et al., 2021a).
+These detailed explanations are crucial for helping developers quickly analyze and locate vulnerability issues, providing essential guidance and reference for effective remediation.
+We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., achieving over a 90% F1 score on the SeVC dataset) and provide detailed explanations.
+We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs in concentrating on vulnerability-prone code, achieving promising results.
+11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
+I Introduction
+A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3, 4, 5].
 These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation.
@@
 Current vulnerability detection techniques mainly include pattern based and deep learning based approaches.
-Pattern based approaches (Yamaguchi, 2017; rou, [n.
-d.]) generally rely on manually defined rules to detect vulnerabilities.
-Deep learning based approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) train the models using existing vulnerability data and various code representation techniques.
+Pattern based approaches [6, 7] generally rely on manually defined rules to detect vulnerabilities.
+Deep learning based approaches [1, 2, 8, 9, 10] train the models using existing vulnerability data and various code representation techniques.
 Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities.
@@
 They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components.
-To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection.
+To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection and explanation.
 This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
-Figure 1. Example of vulnerability detection result with and without explanation.
-The challenges of this fine-tuning process are twofold: obtaining high-quality training data (Zhou et al., 2019; Chen et al., 2023; Zhou et al., 2024; Zhao et al., 2023; Wang et al., 2024) and effectively adapting the models (Xu et al., 2023; Hu et al., 2021; Thapa et al., 2022).
+Figure 1: Example of vulnerability detection result with and without explanation.
+The challenges of this fine-tuning process are twofold: obtaining high-quality training data [1, 11, 12, 13, 14] and effectively adapting the models [15, 16, 17].
 Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial.
@@
 Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
-In this study, we explore enhancing LLMs to detect and explain software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using proprietary LLM models.
+In this study, we explore the enhancement of LLMs for detecting and explaining software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using advanced LLM models.
 The annotated explanatory information includes the location of the vulnerability and a detailed explanation.
-Following this, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance (Fu et al., 2023a).
-Additionally, we investigate the correlation between vulnerability detection and explanation, as well as utilizing LLMs to follow a Chain-of-Thought (CoT) (Wei et al., 2022) strategy and perform a concentrated vulnerability detection on the key code that is prone to software vulnerabilities.
-Furthermore, we address the evaluation of generated vulnerability explanations (Cheshkov et al., 2023) by proposing new evaluation metrics and an automated review method using LLMs.
-Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC (Li et al., 2021c) and DiverseVul (Chen et al., 2023)), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation.
+Following this, we employ Low-Rank Adaptation (LoRA) [16] fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance [18].
+Additionally, we investigate the correlation between the capability of vulnerability detection and explanation, as well as utilize LLMs to follow a Chain-of-Thought (CoT) [19] strategy for concentrated vulnerability detection in key code snippets prone to software vulnerabilities.
+Furthermore, we address the evaluation of generated vulnerability explanations [20] by proposing new evaluation metrics and an automated review method using LLMs.
+Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC [21] and DiverseVul [11]), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation.
 We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
@@
 Paper Organization.
-Section 2 summarizes the related work.
-Section 3 presents the methodology of our study.
-Section 4 discusses the results to our research questions.
-Section 5 discusses the implications of our study.
-Section 6 discusses the threats to validity.
-Section 7 concludes the paper.
-2. Related Work
+Section II summarizes the related work.
+Section III presents the methodology of our study.
+Section IV discusses the results to our research questions.
+Section V discusses the implications of our study.
+Section VI discusses the threats to validity.
+Section VII concludes the paper.
+II Related Work
 In this section, we summarize the related work on vulnerability detection and explanation, respectively.
-2.1. Vulnerability Detection
-Prior studies proposed a series of deep learning approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) to detect the vulnerabilities.
+II-A Vulnerability Detection
+Prior studies proposed a series of deep learning approaches [1, 2, 8, 9, 10] to detect vulnerabilities.
 These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
-Apart from deep learning approaches, recent advancements in large language models (LLMs) have significantly influenced the field of software vulnerability detection.
-Advanced methodologies such as zero-shot prompting (Zhou et al., 2024; Fu et al., 2023b), in-context learning (Dong et al., 2022; Brown et al., 2020a), and fine-tuning (Hu et al., 2021; Xu et al., 2023; Chen et al., 2023) have been employed in vulnerability detection.
+Apart from deep learning approaches, recent advancements in LLMs have significantly influenced the field of software vulnerability detection.
+Advanced methodologies such as zero-shot prompting [12, 22], in-context learning [23, 24], and fine-tuning [16, 15, 11] have been employed in vulnerability detection.
 Cheshkov et al.
-(Cheshkov et al., 2023) evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings.
+[20] evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings.
 Gao et al.
-(Gao et al., 2023) proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods.
+[25] proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods.
 Nong et al.
-(Nong et al., 2024) focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting.
+[26] focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting.
 By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy.
 Sun et al.
-(Sun et al., 2024) further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities.
+[27] further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities.
 They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance.
 Yusuf et al.
-(Yusuf and Jiang, 2024) discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages.
+[28] discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages.
 Steenhoek et al.
-(Steenhoek et al., 2024) surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
+[29] surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
 These studies underscore the challenges and opportunities of applying LLMs to vulnerability detection.
 Unlike prior methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
-2.2. Vulnerability Explanation
+II-B Vulnerability Explanation
 Although significant research progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these techniques for vulnerability explanation remains a challenging issue.
 Currently, only a limited number of studies focus on the explanatory capabilities of deep learning based models regarding vulnerabilities.
-The VulDeeLocator  (Li et al., 2021b) enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer.
-In a similar vein, both IVDetect (Li et al., 2021a) and LineVul (Fu and Tantithamthavorn, 2022) develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements.
-VELVET (Ding et al., 2022) prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks.
-LineVD (Hin et al., 2022) addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning.
-VulTeller (Zhang et al., 2023) emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities.
-VulExplainer  (Cheng et al., 2024) is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks.
-Coca (Cao et al., 2024) is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
+The VulDeeLocator  [30] enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer.
+In a similar vein, both IVDetect [2] and LineVul [31] develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements.
+VELVET [32] prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks.
+LineVD [33] addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning.
+VulTeller [34] emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities.
+VulExplainer  [35] is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks.
+Coca [36] is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
 The semantic understanding of neural network offers the potential for vulnerability explanation models tailored to software development.
@@
 Our work explores the feasibility of fine-tuning specialized LLMs for vulnerability detection and explanation, addressing a crucial gap in this research domain.
-3. Methodology
-Figure 2. Overview of our study.
-To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to identify and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for vulnerability detection and explanation.
+III Methodology
+Figure 2: Overview of our study.
+To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for both vulnerability detection and explanation.
 Figure 2 presents an overview of our framework, namely, LLMVulExp.
@@
 In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important.
-In this paper, we conduct the study on two datasets: (1) SeVC (Li et al., 2021c), which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul (Chen et al., 2023), which covers 295 real open-source projects and 150 CWE types.
-Table 1. Statistics of the studied datasets.
+In this paper, we conduct the study on two datasets: (1) SeVC [21], which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul [11], which covers 295 real open-source projects and 150 CWE types.
+TABLE I: Statistics of the studied datasets.
 Dataset Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
@@
 DiverseVul 18,945 9,161 10 Multi-Type(CWE)
-Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones.
+Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones.
 The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression.
 The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
-DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 Common Weakness Enumerations (CWEs).
+DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 CWEs.
 DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise.
 Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types.
-We select the top ten CWE types to construct our fine-tuning dataset, enabling us to conduct multi-label classification detection
+We select the top ten most frequent CWE types to construct a fine-tuning dataset, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
 We deduplicate the vulnerability samples using a hash method.
-Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset.
+Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset, aiming to reduce training overhead and avoid model bias.
 The dataset details are shown in Table 1.
 We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
-❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets primarily include information within the scope of source code, vulnerability labels, CWE types, and commit messages.
-However, there is still a lack of detailed explanations of the vulnerability logic within the source code, which poses challenges for vulnerability detection techniques to provide a corresponding explanation regarding the detection results.
-Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, involving high labor and time costs.
-To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using open-source LLMs.
-This method leverages the contextual learning and instruction-following capabilities of LLMs, using prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation.
-Prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code localization, and specific explanation.
-Combining instruction-based prompt templates and effective annotated examples to stimulate the model’s contextual learning capabilities ensures the effectiveness of vulnerability explanation annotation.
-In this paper, we use GPT-3.5 (Brown et al., 2020b) by invoking the API provided by OpenAI (ope, 2024) to implement the annotation process.
+❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets predominantly encompass information such as source code, vulnerability labels, CWE types, and commit messages.
+However, they lack detailed explanations of the vulnerability logic within the source code, presenting a significant challenge for vulnerability detection techniques to provide corresponding explanations for the detection results.
+Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, which incurs high labor and time costs.
+To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using LLMs.
+This method leverages the contextual learning and instruction-following capabilities of LLMs, utilizing prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation.
+The prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code location, and specific explanation.
+By combining instruction-based prompt templates with well-annotated examples, we stimulate the model’s contextual learning capabilities, ensuring the effectiveness of vulnerability explanation annotation.
+In this paper, we use GPT-3.5 [37], accessed via the API provided by OpenAI [38], to implement the annotation process.
 To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively.
 This effort fills the current gap in vulnerability explanation data.
-The specific annotation process can be found in Section 4.1.
-❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data provides a large-scale dataset for vulnerability detection and explanation, addressing the data bottleneck issues in the fine-tuning process.
-To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune general-purpose LLMs to enable them to detect and explain specific types of vulnerabilities in real code.
-We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized output.
-To reduce the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021), significantly reducing time and space costs.
-❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations makes effective evaluation of LLM-generated vulnerability explanations an important issue.
-Similar to the challenges faced in annotation, manual evaluation also requires significant human and time resources.
-Furthermore, the effectiveness of vulnerability explanations involves multiple dimensions to provide useful assistance to developers in real development scenarios.
-To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and constructiveness, combined with prompt engineering to develop an automated LLM evaluation method.
-Through expert manual verification, we validate the effectiveness of the specialized vulnerability explanation model’s generated quality and the feasibility of the LLM-based automated evaluation scheme.
-3.1. Vulnerability Interpretation Enhancement Prompting
-Despite the strong code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning, especially in tasks that require a deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities.
+❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data creates a large-scale dataset for vulnerability detection and explanation, addressing data bottleneck issues in the fine-tuning process.
+To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune the general LLMs to enable them to detect and explain specific types of vulnerabilities in real code.
+We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized outputs.
+To minimize the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA [16], significantly reducing time and space costs.
+❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations highlights the need for effective evaluation methods for LLM-generated explanations.
+Similar to the challenges in annotation, manual evaluation demands significant human and time resources.
+Additionally, the effectiveness of vulnerability explanations must be assessed across multiple dimensions to ensure they provide valuable assistance to developers in real-world scenarios.
+To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and actionability.
+By leveraging prompt engineering, we develop an automated LLM evaluation method.
+Expert manual verification is employed to validate both the quality of the specialized vulnerability explanation model’s outputs and the feasibility of the LLM-based automated evaluation scheme.
+III-A Vulnerability Interpretation Enhancement Prompting
+Despite the robust code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning tasks that require deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities.
 These challenges manifest as insufficient detection accuracy and vague vulnerability analysis.
-To guide LLMs towards better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques and the contextual learning capabilities of LLMs.
-By combining these with prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
-Figure 3 shows the design of our prompt templates.
-The prompt templates consist of four main components: task description, specific instructions, generation examples, and sample inputs.
-(1) Task description: provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format, helping the LLM understand task requirements and grasp the background knowledge of the vulnerabilities.
+To guide LLMs toward better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs.
+By leveraging prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
+Figure 3 illustrates the design of our prompt templates.
+These templates consist of four main components: task description, specific instructions, generation examples, and sample inputs.
+(1) Task description: Provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format.
+This helps the LLM understand task requirements and grasp the background knowledge of the vulnerabilities.
 (2) Specific instructions: Include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output.
@@
 (3) Generation examples: Provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals.
-(4) Sample input: is used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
-Figure 3. Prompting template of Automated Vulnerability Interpretation Labelling.
-Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, leading to well-performing fine-tuned specialized vulnerability detection and explanation models.
-This confirms the effectiveness of our prompt design approach.
-3.2. Key Code Extraction Based on Chain-of-Thought (CoT)
+(4) Sample input: Used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
+Figure 3: Prompting template of Automated Vulnerability Interpretation Labelling.
+Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, resulting in well-performing fine-tuned specialized vulnerability detection and explanation models.
+III-B Key Code Extraction Based on Chain-of-Thought (CoT)
 One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets.
-Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs, but these methods are difficult to apply in diverse vulnerability detection scenarios.
+Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs.
+However, these methods are often difficult to apply in diverse vulnerability detection scenarios.
 To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction.
 This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code.
-The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities in the form of a thinking chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
-Figure 4 illustrates the prompt template for key statement extraction designed for LLMs.
+The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities through a thought chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
 The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on.
 After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT.
-This guides the model in reference to the key statements step by step to complete the vulnerability detection, localization, and explanation.
-Figure 4. Prompting template of Key Code Extraction.
-4. Results
+This approach guides the model step by step to reference the key statements to complete the vulnerability detection, localization, and explanation.
+IV Results
 In this section, we discuss the results by proposing and answering the following research questions:
@@
   • RQ4: How does the key code extraction impact detection performance?
-4.1. RQ1: How effective are LLMs in detecting software vulnerabilities?
-In this RQ, we discuss the detection performance of fine-tuned, vulnerability-specialized LLMs in various scenarios.
-4.1.1. Experimental Setup.
-Dataset. As discussed in Section 3, we select the SeVC (Li et al., 2021c) dataset and the DiverseVul dataset (Chen et al., 2023). assessments involving ten CWE types of vulnerabilities.
+IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
+Accurate detection is the foundation for correct vulnerability explanation.
+In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
+IV-A1 Experimental Setup
+Dataset.
+As discussed in Section III, we select the SeVC [21] dataset and the DiverseVul dataset [11].
+Based on the characteristics of the two datasets, we constructed three different vulnerability detection tasks to evaluate the detection performance of LLMVulExp: binary classification vulnerability detection (SeVC), coarse-grained multi-class vulnerability detection (SeVC), and CWE type-based multi-label vulnerability detection (DiverseVul).
 Backbone LLMs.
-We use the Codellama (Roziere et al., 2023) and Llama3 as our base models for fine-tuning.
-CodeLlama is initialized with the weights of Llama2 (Touvron et al., 2023) and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities.
-Considering the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct (cod, 2023) as our primary model and used the 7B version and the newly released Llama3-8B-Instruct for comparison (refer to the discussion section).
-Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability models, we design four sub-research questions:
+We use the Codellama [39] and Llama3 as our backbone LLMs for fine-tuning.
+CodeLlama is initialized with the weights of Llama2 [40] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities.
+The Llama-family LLMs have been widely used by prior studies related to LLMs for software engineering [41, 42, 43].
+In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [44] as our primary model for the experiments discussed in this chapter.
+For comparison purposes, we also considered the 7B version and the newly released Llama3-8B-Instruct in the discussion section.
+Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design four sub-research questions:
   • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
@@
 To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics.
-For the binary classification scenarios in RQ1.1 and RQ1.2, we directly use these three metrics.
-For the multi-class classification in RQ1.3, we use Weighted-F1 and Macro-F1.
-For the multi-label classification in RQ1.4, we use Micro-F1 and Macro-F1.
-To test the detection accuracy, we selected CodeT5 (Wang et al., 2021) and CodeBERT (Feng et al., 2020) as baselines and fine-tuned them for classification tasks by adding a linear classification layer.
+For the binary classification scenario in RQ1.1, we directly use these three metrics.
+For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1.
+For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1.
+To test the detection accuracy, we selected CodeT5 [45] and CodeBERT [46] as baselines, which have been commonly used by prior studies in software engineering [47, 48].
+In particular, we fine-tune them for classification tasks by adding a linear classification layer.
 Implementation Details.
-We implement our approach using the Transformers (Thapa et al., 2022) and PEFT (Mangrulkar et al., 2022) libraries with the PyTorch platform.
-All experiments are conducted on NVIDIA A100-SXM4-80GB GPU platforms, with the token length limit set to 2048.
+We implement our approach using the Transformers [17] and PEFT [49] libraries with the PyTorch platform.
+All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs platforms, with the token length limit set to 2048.
 We use the AdamW optimizer and train the models for 3 epochs.
-For Low-rank Adaptation (LoRA) configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
-4.1.2. RQ1.1: Detection with specified vulnerability type using a dedicated model.
+For LoRA configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
+IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model
 To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type.
-Specifically, we indicate the type of vulnerability we are concerned with in the training and inference prompts to help it narrow down the scope of detecting vulnerabilities.
-During training, we use the explanations with vulnerability location information and specific details annotated by GPT as the target output for the model to perform detection and explanation tasks.
-The inputs for the model during training and inference are code snippets of the given examples.
-During inference, we develop a rule to analyze the generated result to determine whether it indicates a vulnerability or not.
-Specifically, we consider a result as non-vulnerable if it contains the sentences following a fixed pattern (e.g., “There is no security issues” and “detected in the given code snippet.”).
-Otherwise, we consider the result indicates a vulnerability.
-We randomly selected 100 vulnerable results and 100 non-vulnerable results to examine the correctness of such rule-based analysis.
-We find that all of the vulnerable and non-vulnerable results can be correctly classified by this rule.
-Table 2. Performances of Single-Type Detection on SeVC(RQ1.1).
+The sample size of each type is shown in Table II.
+In the training and inference prompts, we specify the type of vulnerability to help narrow the scope of detection.
+During training, we use explanations annotated by GPT-3.5, which include vulnerability location information and specific details, as the target output for the model to perform detection and explanation tasks.
+The inputs for the model during both training and inference are code snippets from the given examples.
+As a generative model, we conduct a binary classification task based on the semantics of the generated text.
+For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it outputs a predefined fixed pattern (e.g., “There are no security issues”).
+TABLE II: Performances Comparison of fine-tuning on SeVC(RQ1.1).
+Metric API Arith. Pointer Array Average
+#Samples 20,294 5,968 38,040 16,680 80,982
+Precision (LLMVulExp) 91.6% 90.3% 93.7% 95.3% 92.7%
+Precision (CodeLlama) 61.8% 64.5% 70.2% 66.2% 65.7%
+Recall (LLMVulExp) 94.5% 93.6% 91.2% 89.7% 92.3%
+Recall (CodeLlama) 26.3% 30.4% 50.2% 36.6% 35.9%
+F1 (LLMVulExp) 93.0% 91.9% 92.4% 92.4% 92.4%
+F1 (CodeLlama) 36.9% 41.4% 58.6% 47.2% 46.0%
+TABLE III: Performances of Single-Type Detection on SeVC(RQ1.1).
 Precision Recall F1
@@
 Experimental Results.
-The experimental results are shown in Table 2.
-The fine-tuned models on four types of vulnerabilities have exceeded 90% performances on three metrics, which indicates that the specialized models for a single type of vulnerability can achieve good results.
-On one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type.
+We first compared the impact of fine-tuning on the detection accuracy of CodeLlama.
+As shown in Table II, CodeLlama-13B-Instruct lacks precise vulnerability identification capabilities.
+Fine-tuning significantly improves the detection accuracy.
+Therefore, in subsequent tasks, we no longer compare with the original CodeLlama.
+From Table III we can observe that our generative model still achieves comparable performance to other fine-tuned classification models.
+On the one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type.
 On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs.
 If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
-4.1.3. RQ1.2: Detection with specified vulnerability type using an all-in-one model.
-To explore the impact of training with data from other types of vulnerabilities on the detection effect of the target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning on four types of vulnerabilities and performing individual detection for each type.
-Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection, and during inference, we also provide the vulnerability type information for each example.
-The difference between RQ1.1 and RQ1.2 is that, we fine-tune a dedicated model for each type of vulnerability in RQ1.1.
-In RQ1.2, we fine-tune an all-in-one model and use the prompts to distinguish the type of vulnerability in interest.
-Table 3. Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
+IV-A3 RQ1.2: Detection with specified vulnerability type using an all-in-one model
+To explore the impact of training with data from different types of vulnerabilities on the detection effectiveness for target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning a model on four types of vulnerabilities and then performing individual detection for each type.
+Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection.
+During inference, we also provide the vulnerability type information for each example.
+The difference between RQ1.1 and RQ1.2 is that in RQ1.1, we fine-tune a dedicated model for each type of vulnerability, whereas in RQ1.2, we fine-tune an all-in-one model and use prompts to distinguish the type of vulnerability of interest.
+TABLE IV: Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
 Accuracy Precision Recall F1
@@
 Experimental Results.
-We present the results of the fine-tuned all-in-one model in Table 3.
+We present the results of the fine-tuned all-in-one model in Table IV.
 In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities.
@@
 Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
-4.1.4. RQ1.3: Detection with an identification of the vulnerability type.
-In real software development environments, the vulnerability risks faced are often diverse.
-It may not always be possible to preemptively sense the potential types of vulnerabilities in the samples to be detected.
-This requires the large model to correctly identify the type of vulnerabilities present from multiple types and further provide specific explanations.
-To test the accuracy of a unified model in identifying and classifying code with multiple types of vulnerabilities, we conduct a multi-class vulnerability detection fine-tuning on SeVC.
-Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including none-vulnerable and the four types of vulnerabilities in SeVC.
-In the training and inference prompts, we no longer specify the type of vulnerability, but instead turn to a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
-Table 4. Performances of Multi-Type Detection on SeVC(RQ1.3).
-Precision Recall F1 Score
+IV-A4 RQ1.3: Detection with an identification of the vulnerability type
+In real software development environments, the vulnerability risks faced are often diverse, making it difficult to preemptively sense the potential types of vulnerabilities in the samples to be detected.
+This requires the large model to accurately identify and classify multiple types of vulnerabilities and provide specific explanations.
+To test the accuracy of a unified model in this context, we conduct a multi-class vulnerability detection fine-tuning on SeVC with all samples.
+Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including non-vulnerable and the four types of vulnerabilities in SeVC.
+In the training and inference prompts, we no longer specify the type of vulnerability.
+Instead, we adopt a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
+TABLE V: Performances of Multi-Type Detection on SeVC(RQ1.3).
+Precision Recall F1
 Type Ours CodeT5 CodeBert Ours CodeT5 CodeBert Ours CodeT5 CodeBert
@@
 Experimental Results.
-We present the metrics for each type and the confusion matrix for the multi-class vulnerability detection task in Table 4.
+We present the metrics for each type for the multi-class vulnerability detection task in Table V.
 Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code.
@@
 The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
-4.1.5. RQ1.4: Identification of the vulnerability type on the datasets with more vulnerability types.
-In real development environments, the security risks of vulnerabilities come from a greater variety of project types and more complex code structures, often involving a wider range of vulnerability models.
-To more closely resemble real development conditions, we conduct a multi-label classification task of the top 10 types of CWE on the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels).
-Specifically, we modify the task description to be based on CWE-type vulnerability detection and explanation tasks, and add the ‘[CWE]’ tag for generating a list of CWEs related to the target code.
-At the same time, we incorporate the CWE descriptions as part of the output explanation information to help the model further understand the meaning of each CWE type of vulnerability and to use the CWE descriptions for vulnerability analysis and identification.
-Table 5. Performances of Multi-Label CWE Detection on DiverseVul(RQ1.4).
-Metric CWE ID Overall
-787 125 119 20 703 416 476 190 200 399 Micro Macro
-Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 73.3%
-Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 73.4%
-F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 73.2%
+IV-A5 RQ1.4: Identification of the vulnerability type in datasets with more vulnerability types
+In real development environments, security risks stem from a wide variety of project types and complex code structures, often involving a broader range of vulnerability models.
+To better resemble real development conditions, we conduct a multi-label classification task on the top 10 types of CWE in the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels).
+Specifically, we modify the task description to focus on CWE-type vulnerability detection and explanation tasks, adding a ‘[CWE]’ tag to generate a list of CWEs related to the target code.
+Additionally, we incorporate CWE descriptions as part of the output explanation to help the model better understand the meaning of each CWE type and utilize these descriptions for vulnerability analysis and identification.
+TABLE VI: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.4).
+Metric CWE ID Overall (Micro)
+787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
+Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 70.8% 65.3%
+Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 63.1% 61.9%
+F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 67.0% 63.5%
 Experimental Results.
-The experimental results are presented in Table 5.
-From the results, we find that after fine-tuning, the specialized vulnerability models have achieved good detection accuracy for the 10 types of CWE, demonstrating an ability to grasp the pattern information for each CWE type.
+The experimental results are presented in Table VI.
+From the results, we find that after fine-tuning, the specialized vulnerability models achieve good detection accuracy for the 10 types of CWE, demonstrating their ability to grasp the pattern information for each CWE type.
 Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories.
-Each vulnerability type has a more fine-grained explicit definition, and it also avoids the potential category association issues that might exist in SeVC, thus resulting in better performance.
-This validates the feasibility of research into LLMs for vulnerability detection and explanation that are oriented toward real software development environments.
+Each vulnerability type has a more fine-grained and explicit definition, avoiding potential category association issues in SeVC and resulting in better performance.
+This validates the feasibility of using LLMs for vulnerability detection and explanation in real software development environments, which is illustrated in the following explanation example.
+1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.
+[cwe] This function is related to [’CWE-476’].
+CWE-476 NULL Pointer Dereference: A NULL pointer dereference occurs when the application dereferences a pointer that it expects to be valid, but is NULL, typically causing a crash or exit.
+[location] if (!wasm_->onNewConnection_)
+[explanation] The function is vulnerable to a NULL pointer dereference.
+It checks if ’wasm_->onNewConnection_’ is NULL but does not handle the case when it is NULL.
+If ’wasm_->onNewConnection_’ is NULL, the function continues its execution, which could lead to a NULL pointer dereference and cause a crash or exit.
+The vulnerability is related to CWE-476: NULL Pointer Dereference.
+TABLE VII: Example of Generated Explanation on DiverseVul.
 Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
-4.2. RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
+IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
 In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities.
-Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanation.
+Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanations.
 The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content.
 Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
-Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics, Accuracy, Clarity, and Constructiveness, as the evaluation criteria for vulnerability explanation.
-We use manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
-4.2.1. Experimental Setup.
-In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT.
-We randomly sample the data based on 95% confidence level and 5% confidence interval (Boslaugh and Watters, 2008) for the fine-tuned model’s correctly classified instances, obtaining 384 and 224 samples from SeVC and DiverseVul, respectively.
-Our evaluation includes two parts, (1) Manual review: two authors of this paper independently conduct manual checks on the samples, scoring them according to the criteria.
-In cases of disagreement, discussions are held until a consensus.
-The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement.
+Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics: Accuracy, Clarity, and Actionability, as the evaluation criteria for vulnerability explanations.
+We use both manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
+IV-B1 Experimental Setup
+In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT-3.5.
+Due to the large scale of the dataset we use and the high evaluation cost, we randomly sample the data based on a 95% confidence level and a 5% confidence interval [50] following prior studies [51, 52, 53] for the fine-tuned model’s correctly classified instances.
+We obtain 384 and 224 samples from SeVC and DiverseVul, respectively.
+This sampling method ensures the statistical validity and representativeness of our evaluation.
+Our evaluation includes two parts: (1) Manual review: Two authors of this paper, both of whom have extensive research experience in software vulnerabilities and software development, independently conduct manual checks on the samples, scoring them according to the criteria.
+In cases of disagreement, discussions are held until a consensus is reached.
+The value of Cohen’s Kappa [54] in this process is 0.76, indicating substantial agreement.
+This rigorous manual review process ensures the reliability and accuracy of our evaluation.
 (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion.
@@
   • Clarity: The information should be presented in a clear, understandable manner, and structured in a way that facilitates easy comprehension by software developers.
-  • Constructiveness: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
-4.2.2. Experimental Results.
-Table 6. Distribution (%) of Vulnerability Explanation Review Results (RQ2).
+  • Actionability: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
+IV-B2 Experimental Results
+We present the results of our dual evaluation process in Table VIII, the suffix Gen.
+refers to the explanation generated by LLMVulExp and Ann.
+refers to the explanation annotated by GPT-3.5.
+All-Pos.
+refers to the results that meet the requirements of all of the three metrics.
+Combining the results from manual and LLM evaluations, we have made the following observations:
+(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen.
+for both the manual review and automated review) in the explanation.
+They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
+(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively.
+The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
+(3) Actionability of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
+(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann.
+and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data.
+It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
+(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review.
+For SeVC and DiverseVul, the results of All-Pos.
+comparing manual review and automated review are 76.0% vs.
+74.0% and 80.1% vs.
+71.9%, respectively.
+This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text.
+In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
+These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation.
+The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations.
+The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities.
+The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis.
+Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
+TABLE VIII: Distribution (%) of Vulnerability Explanation Review Results (RQ2).
 Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
@@
 All-Pos. 72.8 74.0 67.9 71.9
-We present the results of our dual evaluation process in Table 6, the suffix Gen.
-refers to the explanation generated by LLMVulExp and Ann.
-refers to the explanation annotated by GPT.
-All-Pos.
-refers to the results that meet requirements of all of the three metrics.
-Combining the results from manual and LLM evaluations, we have made the following observations:
-(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen.
-for both the manual review and automated review) in the explanation.
-They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
-(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively.
-The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
-(3) Constructiveness of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
-(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann.
-and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data.
-It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
-(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review.
-For SeVC and DiverseVul, the results of All-Pos.
-comparing manual review and automated review are 76.0% vs.
-74.0% and 80.1% and 71.9%, respectively.
-This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text.
-In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
-These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation.
-The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations.
-The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities.
-The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis.
-Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
-Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Constructiveness.
+1int i; 2for(i = 0; i < 1; i++) 3twoIntsStruct * dataBuffer = (twoIntsStruct *) ALLOCA(100 * sizeof(twoIntsStruct)); 4size_t i; 5for (i = 0; i < 100; i++) 6 dataBuffer[i].intOne = 1; 7 dataBuffer[i].intTwo = 1; 8data = dataBuffer; 9printStructLine(&data[0]); 10 11void printStructLine(const twoIntsStruct * structTwoIntsStruct) { 12 printf($d\n$", structTwoIntsStruct->intOne, structTwoIntsStruct->intTwo); 13} 14delete [] data;
+{mdframed}[style=mystyle] [type] pointer
+[location] The line ”delete [] data” has a pointer-related issue.
+[explanation]
+(Analysis:) The issue in this code is related to the ”data” pointer, which is being improperly freed using the ”delete” operator.
+The ”data” pointer is assigned the address of a memory block allocated on the stack using the ”ALLOCA” function, which allocates memory on the stack instead of the heap.
+The ”delete” operator is used to free memory allocated on the heap using the ”new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
+(Suggestion:) To fix this issue, the program should use the ”free” function to free memory allocated on the stack, or use the ”new” operator to allocate memory on the heap instead of the ”ALLOCA” function.
+Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
+TABLE IX: Example of Generated Explanation on SeVC.
+IV-B3 Case Study
+To more intuitively illustrate the importance of evaluating the quality of vulnerability explanations, we analyze an example generated by our model.
+As shown in Figure IV-B2, this vulnerability code is sourced from SeVC and pertains to pointer-type vulnerabilities, demonstrating incorrect memory deallocation in C language using the delete operator.
+The explanation adheres to our output paradigm, generating vulnerability detection, location, and detailed information.
+The provided explanation correctly identifies the core issue: the data pointer is allocated on the stack using the ALLOCA function but is improperly freed using the delete operator.
+The explanation points out that the delete operator should only be used to free memory allocated on the heap via the new operator.
+Using it to free stack-allocated memory can lead to undefined behavior and potential crashes, meeting the accuracy requirement.
+Additionally, the explanation offers recommendations for addressing this vulnerability, such as using the free function to release stack-allocated memory or using the new operator for heap allocation, and ensuring pointer safety.
+This satisfies the actionability requirement.
+The explanation is clear, comprehensive, logically coherent, and easy to read, meeting the clarity requirement.
+During the review, we also observe some instances with insufficient quality of explanations.
+Issues include failure to identify the main vulnerability, vague location, verbose expression, and lack of remediation suggestions.
+Hallucinations in explanation generation are relatively rare but primarily involved memorized descriptions of similar code samples from the training set and location confusion.
+Overall, the fine-tuned model demonstrates a certain capability for generating vulnerability explanations.
+Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Actionability.
 We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
-4.3. RQ3: How do explanations affect the results of vulnerability detection?
-By considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1.
-This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs, compared to fine-tuning specialized detection models?
-In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection (i.e., comparing the performance of vulnerability detection between fine-tuned LLMs with and without the explanatory information).
+IV-C RQ3: How do explanations affect the results of vulnerability detection?
+Considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1.
+This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs compared to fine-tuning specialized detection models?
+In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory information.
 Experimental Setup.
-We conduct our research in the form of ablation studies under the settings of multi-type vulnerability detection scenario in SeVC (RQ1.3) and multi-label CWE detection scenario in DiverseVul (RQ1.4).
+We conduct our research as ablation studies under the settings of multi-type vulnerability detection scenarios in SeVC (RQ1.3) and multi-label CWE detection scenarios in DiverseVul (RQ1.4).
 We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt.
@@
 We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
-Figure 5. Performance Comparison between Explainer and Detector.
+Figure 4: Performance Comparison between Explainer and Detector.
 Experimental Results.
-We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 5, the prefix of “Mi” and “Ma” refers to Micro and Macro.
-Based on the experimental results, incorporating the vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance.
-For example, the weightd precision on SeVC for Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% and 74.8% on DiverseVul.
+We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 4.
+The prefix of “Mi” and “Ma” refers to Micro and Macro, respectively.
+Based on the experimental results, incorporating vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance.
+For example, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% for both on DiverseVul.
 In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
 These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities.
-The model’s detection capability still remains with an additional focus on explanation capabilities.
+The model’s detection capability remains intact with an additional focus on explanation capabilities.
 Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns.
-It suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance.
+This suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance.
 Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
 Summary of RQ3: Integrating vulnerability explanation into the fine-tuning process does not compromise detection capabilities and may even lead to improved performance for certain vulnerability scenarios.
-4.4. RQ4: How does the key code extraction impact detection performance?
-In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to the vulnerability and further examine the code with a specific focus.
+IV-D RQ4: How does the key code extraction impact detection performance?
+In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to vulnerability and further examine the code with a specific focus.
 To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
 Experimental Setup.
-Similar to the process of annotation discussed in Section 3, we use GPT-3.5 to extract the key code.
-The prompt templates are illustrated in Figure 4.
+Similar to the process of annotation discussed in Section III, we use GPT-3.5 to extract the key code.
 To prevent label leakage, the code vulnerability type tags were not visible during the annotation process.
@@
 During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
-Table 7. F1 Scores of Key Code Extraction on SeVC (RQ4)
+TABLE X: F1 Scores of Key Code Extraction on SeVC (RQ4).
 Type Single-Type(RQ1.1) Multi-Type(RQ1.3)
@@
 Experimental Results.
-We present the experimental results of utilizing Key Code Extraction on SeVC in Table 7 and DiverseVul in Figure 6.
+We present the experimental results of utilizing Key Code Extraction on SeVC in Table X and DiverseVul in Figure 5.
 We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type).
@@
 This makes it easier for the model to learn and detect patterns associated with vulnerabilities.
-Figure 6. Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent for the CWE IDs.
-Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., 64.4% -¿ 83.8% for Multi-Type detection on SeVC).
-5. Discussion
+This method does not rely on manual feature extraction, making it broadly applicable.
+Our experimental results preliminarily demonstrate the feasibility of key code extraction through LLMs.
+Figure 5: Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent the CWE IDs.
+Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
+V Discussion
 In this section, we discuss the implications of our study.
-Implication 1: Implications for Broader Vulnerability-Related Tasks.
-Our study reveals that integrating the task of vulnerability explanation with detection does not compromise the model’s performance.
-This finding suggests we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs.
+Implication 1: Broader Applications for Vulnerability-Related Tasks.
+Our study reveals that integrating vulnerability explanation with detection does not compromise the model’s performance.
+This finding suggests that we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs.
 By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights.
@@
 In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments.
-To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs (i.e., CodeLlama-7B-Instruct and Llama3-8B-Instruct) to conduct the experiments following the setting of RQ1.3.
-Figure 7 shows the results of Multi-Type vulnerability detection using LLMs with different sizes.
-We find that LLMVulExp can also achieve effective performance using the additional LLMs.
-Particularly, the performance between CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct.
-Overall, LLMVulExp presents its effectiveness using different LLMs with various sizes.
-Figure 7. Performance of Multi-Type vulnerability detection using LLMs with different sizes.
+To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, to conduct the experiments following the setting of RQ1.3.
+Figure 6 shows the results of multi-type vulnerability detection using LLMs of different sizes.
+We find that LLMVulExp can achieve effective performance with these additional LLMs.
+Notably, the performance of CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct.
+Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
+Figure 6: Performance of Multi-Type vulnerability detection using LLMs with different sizes.
 Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks.
@@
 This hybrid approach could balance the need for detailed annotations and the practical constraints of manual labeling.
-6. Threats to Validity
+VI Threats to Validity
 Construct Validity.
-We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Constructiveness.
-These metrics characterize three aspects of effective vulnerability explanations but may not comprehensively cover all features of vulnerability explanations.
+We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Actionability.
+While these metrics effectively characterize important aspects of vulnerability explanations, they may not comprehensively cover all features.
 More comprehensive and refined evaluation metrics for vulnerability explanations require further research.
-In this paper, we employed both manual review and automated review using LLMs.
-To mitigate the bias, two authors independently annotated the explanations and resolved any discrepancies until a consensus.
-The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement.
+To mitigate potential bias, we employed both manual and automated reviews using LLMs.
+Two authors independently annotated the explanations and resolved any discrepancies until a consensus was reached.
+The value of Cohen’s Kappa [54] in this process was 0.76, indicating substantial agreement.
+This rigorous approach helps ensure the reliability and validity of our evaluation.
 Internal Validity.
@@
 The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results.
-However, we conducted at least two or more trials to validate the stability of our conclusions for each experimental setup.
-The number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited.
-Additionally, due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version.
-Although these limitations may have resulted in alternative findings, they do not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
+However, we conducted multiple trials to validate the stability of our conclusions for each experimental setup.
+Although the number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited, and due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version, we ensured that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
 External Validity.
@@
 However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models.
-We conducted experiments using two differently configured open-source project-based vulnerability datasets.
-While these datasets contain a lot of vulnerable code, they differ from real-world development environments.
-Therefore, our findings may not apply to other datasets or industrial settings.
-7. Conclusion
-In this paper, we propose LLMVulExp, a framework that aims that detecting and explaining vulnerabilities using LLMs.
+Additionally, we conducted experiments using two differently configured open-source project-based vulnerability datasets.
+Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments.
+Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments.
+Nonetheless, our careful selection and thorough evaluation process aimed to ensure that our findings remain relevant and insightful within the scope of our study.
+VII Conclusion
+In this paper, we propose LLMVulExp, a framework designed to detect and explain vulnerabilities using LLMs.
 The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security.
 Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation.
-It suggests that a multi-task learning approach could be beneficial, addressing the data volume bottleneck is crucial for practical applications, and the quality of annotations is paramount for the success of fine-tuning frameworks.
-By considering these insights, future work can aim to develop more capable and efficient models that contribute to the field of software security.
+It highlights the importance of addressing the data volume bottleneck for training vulnerability LLMs in software development.
+Furthermore, the quality of annotations is paramount for the success of fine-tuning frameworks.
+Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.

### v3

@@
 Software vulnerabilities pose significant risks to the security and integrity of software systems.
-Prior studies have proposed various approaches to vulnerability detection using deep learning or pre-trained models.
-However, there is still a lack of detailed explanations for understanding vulnerabilities beyond merely detecting their occurrence, which fails to truly help software developers understand and remediate the issues.
-Recently, large language models (LLMs) have demonstrated remarkable capabilities in comprehending complex contexts and generating content, presenting new opportunities for both detecting and explaining software vulnerabilities.
-In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in both detecting and explaining vulnerabilities, and we propose LLMVulExp, a framework that utilizes LLMs for these tasks.
-Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities.
-These detailed explanations are crucial for helping developers quickly analyze and locate vulnerability issues, providing essential guidance and reference for effective remediation.
-We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., achieving over a 90% F1 score on the SeVC dataset) and provide detailed explanations.
-We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs in concentrating on vulnerability-prone code, achieving promising results.
+Although prior studies have explored vulnerability detection using deep learning and pre-trained models, these approaches often fail to provide the detailed explanations necessary for developers to understand and remediate vulnerabilities effectively.
+The advent of large language models (LLMs) has introduced transformative potential due to their advanced generative capabilities and ability to comprehend complex contexts, offering new possibilities for addressing these challenges.
+In this paper, we propose LLMVulExp, an automated framework designed to specialize LLMs for the dual tasks of vulnerability detection and explanation.
+To address the challenges of acquiring high-quality annotated data and injecting domain-specific knowledge, LLMVulExp leverages prompt-based techniques for annotating vulnerability explanations and fine-tunes LLMs using instruction tuning with Low-Rank Adaptation (LoRA), enabling LLMVulExp to detect vulnerability types in code while generating detailed explanations, including the cause, location, and repair suggestions.
+Additionally, we employ a Chain-of-Thought (CoT) based key code extraction strategy to focus LLMs on analyzing vulnerability-prone code, further enhancing detection accuracy and explanatory depth.
+Our experimental results demonstrate that LLMVulExp achieves over a 90% F1 score on the SeVC dataset, effectively combining high detection accuracy with actionable and coherent explanations.
+This study highlights the feasibility of utilizing LLMs for real-world vulnerability detection and explanation tasks, providing critical insights into their adaptation and application in software security.
 11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
 I Introduction
-A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3, 4, 5].
+A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2].
 These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation.
@@
 Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
-In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue.
-LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation.
-These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations.
-However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation.
-They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components.
-To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection and explanation.
-This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
-Figure 1: Example of vulnerability detection result with and without explanation.
-The challenges of this fine-tuning process are twofold: obtaining high-quality training data [1, 11, 12, 13, 14] and effectively adapting the models [15, 16, 17].
-Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial.
-These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws.
-Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation.
-Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
-In this study, we explore the enhancement of LLMs for detecting and explaining software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using advanced LLM models.
-The annotated explanatory information includes the location of the vulnerability and a detailed explanation.
-Following this, we employ Low-Rank Adaptation (LoRA) [16] fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance [18].
-Additionally, we investigate the correlation between the capability of vulnerability detection and explanation, as well as utilize LLMs to follow a Chain-of-Thought (CoT) [19] strategy for concentrated vulnerability detection in key code snippets prone to software vulnerabilities.
-Furthermore, we address the evaluation of generated vulnerability explanations [20] by proposing new evaluation metrics and an automated review method using LLMs.
-Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC [21] and DiverseVul [11]), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation.
-We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
+In the context of enhancing detection methods, the advent of Large Language Models (LLMs) [8, 9, 10] offers a promising avenue.
+With their advanced generative capabilities, LLMs have demonstrated significant potential across a wide range of applications, such as natural language processing and machine translation.
+These models can generate extensive textual content and provide contextually relevant information, making them well-suited for tasks that require detailed explanations.
+However, a considerable gap exists between the current capabilities of LLMs, particularly open-source models, and the specific requirements of vulnerability detection and explanation.
+Without domain-specific knowledge of vulnerabilities, LLMs struggle to effectively detect vulnerabilities and offer accurate explanations, which demand a deep understanding of code structures, security contexts, and the intricate interplay between various software components.
+To bridge this gap, it is essential to enhance LLMs with specialized capabilities for vulnerability detection and explanation.
+Fine-tuning open-source LLMs [11, 12, 13] on task-specific data presents an effective path toward achieving model specialization.
+Figure 1: Illustration of Vulnerability Detection Explanation.
+The main challenges in this specialization fine-tuning lie in two areas: (1) acquiring high-quality vulnerability explanation data [1, 14, 10, 15, 16], and (2) injecting vulnerability-related domain knowledge into LLMs [17].
+The improvement of model specialization depends heavily on high-quality, domain-specific data.
+However, in the field of vulnerability detection, there is a significant lack of annotated explanation data.
+Additionally, annotating such data typically requires costly manual labeling by domain experts, making the collection of large-scale annotated datasets a major challenge for model fine-tuning.
+Additionally, due to the inherent complexity of vulnerability detection and explanation tasks, merely providing task descriptions in prompts is insufficient for generating accurate answers.
+The model requires additional, critical vulnerability domain knowledge to identify and explain vulnerabilities accurately.
+To tackle these challenges, this study introduces LLMVulExp, an innovative automated framework designed to specialize open-source LLMs for software vulnerability tasks.
+Our framework directly addresses the two key challenges identified: the acquisition of high-quality annotated explanatory data and the effective injection of domain-specific vulnerability knowledge into LLMs.
+To address the first challenge, we leverage prompt-based techniques to automatically annotate explanatory information for vulnerability data.
+This annotated data includes both the vulnerability location and a detailed explanation, providing a rich dataset for fine-tuning.
+This approach reduces reliance on costly manual labeling by domain experts, enabling the collection of large-scale, annotated datasets necessary for model specialization.
+In addition, we propose new evaluation metrics for vulnerability explanations and an automated ’LLM-as-a-judge’ evaluation method [18, 19], ensuring a robust and scalable approach to verify the correctness and coherence of the generated explanations.
+For the second challenge, we apply instruction tuning [20] with Low-Rank Adaptation (LoRA) fine-tuning methods [12] to refine the open-source LLMs for injecting sufficient vulnerability patterns and knowledge.
+By focusing on task-specific adaptation, we guide the model to develop a deeper understanding of vulnerability semantics.
+To further enhance the model’s structural understanding of vulnerabilities, we adopt a Chain-of-Thought (CoT) strategy [21], guiding the model to concentrate on key code snippets most prone to vulnerabilities.
 Overall, the contributions of this paper are threefold:
-  • We present a comprehensive and effective workflow for training, inferring, and evaluating LLMs based on open-source vulnerability data.
-  • We pioneer the exploration of the effectiveness of vulnerability explanation models and propose novel evaluation methods and dimensions for these explanations.
-  • We analyze the interrelationship between explanatory and detection capabilities, offering valuable insights for the development of explainable vulnerability detection methodologies.
+  • We introduce a comprehensive and effective workflow for training, inferring, and evaluating LLMs specifically on vulnerability detection and explanation tasks.
+  • We pioneer the exploration of the effectiveness of LLM-based vulnerability explanation generation and propose a tailored evaluation method and metrics for these explanations.
+  • We conduct extensive experiments to evaluate and analyze the feasibility of LLM-based vulnerability explanations, providing valuable insights for the practical application of specialized models in vulnerability detection.
 Paper Organization.
@@
 Section III presents the methodology of our study.
-Section IV discusses the results to our research questions.
+Section IV discusses the results of our research questions.
 Section V discusses the implications of our study.
@@
 II Related Work
-In this section, we summarize the related work on vulnerability detection and explanation, respectively.
+In this section, we review the related work in two key areas: vulnerability detection and explanation.
 II-A Vulnerability Detection
 Prior studies proposed a series of deep learning approaches [1, 2, 5, 6, 7] to detect vulnerabilities.
-These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
-Apart from deep learning approaches, recent advancements in LLMs have significantly influenced the field of software vulnerability detection.
-Advanced methodologies such as zero-shot prompting [12, 22], in-context learning [23, 24], and fine-tuning [16, 15, 11] have been employed in vulnerability detection.
-Cheshkov et al.
-[20] evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings.
-Gao et al.
-[25] proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods.
-Nong et al.
-[26] focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting.
-By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy.
-Sun et al.
-[27] further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities.
-They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance.
-Yusuf et al.
-[28] discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages.
-Steenhoek et al.
-[29] surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
-These studies underscore the challenges and opportunities of applying LLMs to vulnerability detection.
-Unlike prior methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
+These methods have employed labeled vulnerability data to train neural networks, enabling the models to capture semantic features associated with vulnerabilities.
+The advent of LLMs has significantly influenced the field of vulnerability detection.
+Techniques such as zero-shot prompting [10, 9], in-context learning [22, 23], and fine-tuning [12, 11, 14] have been explored to enhance LLM-based vulnerability detection.
+Cheshkov et al.[18] evaluated ChatGPT and GPT-3, highlighting their inability to classify vulnerable code accurately in binary and multi-label settings.
+Gao et al.[24] proposed a benchmark for LLMs in vulnerability detection, demonstrating that with few-shot prompting on simpler datasets, LLMs can perform comparably to deep learning-based methods.
+Nong et al.[25] showed that chain-of-thought prompting, based on the semantic structure of code, improves detection accuracy.
+Sun et al.[26] found that supplementing LLMs with high-quality vulnerability-related knowledge enhances their performance.
+Yusuf et al.[27] observed that natural language instructions boost vulnerability detection across multiple programming languages.
+Steenhoek et al.[28] surveyed eleven LLMs, indicating the limitations of applying LLMs directly to vulnerability detection without fine-tuning.
+These studies underscore the challenges and potential of using LLMs for vulnerability detection.
+Unlike these methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
 II-B Vulnerability Explanation
-Although significant research progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these techniques for vulnerability explanation remains a challenging issue.
-Currently, only a limited number of studies focus on the explanatory capabilities of deep learning based models regarding vulnerabilities.
-The VulDeeLocator  [30] enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer.
-In a similar vein, both IVDetect [2] and LineVul [31] develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements.
-VELVET [32] prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks.
-LineVD [33] addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning.
-VulTeller [34] emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities.
-VulExplainer  [35] is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks.
-Coca [36] is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
-The semantic understanding of neural network offers the potential for vulnerability explanation models tailored to software development.
-However, accurate detection and effective explanation of vulnerabilities using LLMs remain substantial challenges for general-purpose code models.
-Our work explores the feasibility of fine-tuning specialized LLMs for vulnerability detection and explanation, addressing a crucial gap in this research domain.
+Although substantial progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these methods for vulnerability explanation remains a significant challenge.
+Despite the critical importance of reducing reliance on security experts and aiding developers in timely and thorough vulnerability mitigation, this area remains underexplored.
+Only a few studies focus on the explanatory capabilities of deep learning based models for vulnerabilities.
+VulDeeLocator [29] enhances a Bi-LSTM detector with an inner multiplication layer to forecast vulnerable statements.
+IVDetect [2] and LineVul [30] utilize subgraphs or attention weights from trained detectors to identify vulnerabilities.
+VELVET [31] integrates graph-based and sequence-based neural networks to prioritize vulnerable statements.
+LineVD [32] treats statement-level vulnerability detection as a node classification task, combining graph neural networks with a transformer-based model on the Program Dependency Graph(PDG).
+VulTeller [33] focuses on control and taint flows to detect dependencies for localizing vulnerabilities.
+VulExplainer [17] uses graph neural networks to locate fine-grained vulnerability information.
+While neural network-based models offer promising semantic understanding for vulnerability explanation, accurately detecting and explaining vulnerabilities using LLMs remains a substantial challenge for general-purpose code models.
+Our work explores the feasibility of fine-tuning specialized LLMs for both vulnerability detection and explanation, addressing a critical gap in the research literature.
 III Methodology
-Figure 2: Overview of our study.
-To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for both vulnerability detection and explanation.
+Figure 2: Overview of the LLMVulExp Framework for Explainable Vulnerability Detection via LLM Fine-Tuning.
+To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and explain software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized LLMs for both vulnerability detection and explanation tasks.
 Figure 2 presents an overview of our framework, namely, LLMVulExp.
-Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
-❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data.
-In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important.
-In this paper, we conduct the study on two datasets: (1) SeVC [21], which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul [11], which covers 295 real open-source projects and 150 CWE types.
+Specifically, our framework consists of four core stages: ❶ fine-tuning and evaluation vulnerability dataset collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
+❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs requires large quantities of high-quality domain-specific data.
+In the context of vulnerability detection and explanation, it is crucial to ensure the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples for each type.
+In this paper, we conduct the study on two C/C++ function-level vulnerability datasets: (1) SeVC [34], which contains four core vulnerability types and over 50,000 vulnerable code snippets, and (2) DiverseVul [14], which covers 295 real open-source projects and 150 CWE types.
+We select C/C++ as the programming language due to its diverse range of vulnerability types and examples, as well as the substantial dependence of fine-tuning on the availability of large-scale data.
 TABLE I: Statistics of the studied datasets.
@@
 DiverseVul 18,945 9,161 10 Multi-Type(CWE)
-Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones.
-The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression.
-The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
-DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 CWEs.
-DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise.
-Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types.
-We select the top ten most frequent CWE types to construct a fine-tuning dataset, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
-We deduplicate the vulnerability samples using a hash method.
-Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset, aiming to reduce training overhead and avoid model bias.
-The dataset details are shown in Table 1.
-We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
-❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets predominantly encompass information such as source code, vulnerability labels, CWE types, and commit messages.
-However, they lack detailed explanations of the vulnerability logic within the source code, presenting a significant challenge for vulnerability detection techniques to provide corresponding explanations for the detection results.
-Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, which incurs high labor and time costs.
-To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using LLMs.
-This method leverages the contextual learning and instruction-following capabilities of LLMs, utilizing prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation.
-The prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code location, and specific explanation.
-By combining instruction-based prompt templates with well-annotated examples, we stimulate the model’s contextual learning capabilities, ensuring the effectiveness of vulnerability explanation annotation.
-In this paper, we use GPT-3.5 [37], accessed via the API provided by OpenAI [38], to implement the annotation process.
-To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively.
-This effort fills the current gap in vulnerability explanation data.
-❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data creates a large-scale dataset for vulnerability detection and explanation, addressing data bottleneck issues in the fine-tuning process.
-To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune the general LLMs to enable them to detect and explain specific types of vulnerabilities in real code.
-We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized outputs.
-To minimize the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA [16], significantly reducing time and space costs.
-❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations highlights the need for effective evaluation methods for LLM-generated explanations.
-Similar to the challenges in annotation, manual evaluation demands significant human and time resources.
-Additionally, the effectiveness of vulnerability explanations must be assessed across multiple dimensions to ensure they provide valuable assistance to developers in real-world scenarios.
-To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and actionability.
-By leveraging prompt engineering, we develop an automated LLM evaluation method.
-Expert manual verification is employed to validate both the quality of the specialized vulnerability explanation model’s outputs and the feasibility of the LLM-based automated evaluation scheme.
+Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types, comprising 56,395 vulnerable samples and 364,232 non-vulnerable ones.
+The SeVC dataset is categorized into four primary vulnerability classes based on the underlying causes: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression.
+Each category contains a substantial number of samples, making SeVC ideal for fine-tuning and evaluating models designed to detect and explain specific vulnerabilities.
+Therefore, we conduct both binary and multi-class detection tasks on the SeVC dataset.
+DiverseVul is a comprehensive C/C++ vulnerability dataset that includes 18,945 vulnerable functions and 330,492 non-vulnerable functions, extracted from 7,514 commits, covering 150 CWEs.
+As the largest real-world C/C++ vulnerability dataset, DiverseVul is characterized by longer code snippets, a broader range of projects, more diverse vulnerability types, and lower label noise.
+It serves as a benchmark for evaluating the capability of LLMs in handling real-world scenarios with a wider variety of vulnerabilities.
+For our fine-tuning and evaluation target, we selected the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
+To ensure data quality, we deduplicate the vulnerability samples using the SHA-256 [35] hash method and downsample the non-vulnerable samples at a 1:1 ratio to match the vulnerable samples, achieving a balanced dataset, which aims to reduce training overhead and mitigates potential model bias.
+The details of the datasets are shown in Table I.
+We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerable and non-vulnerable samples to conduct our experiments.
+❷ Automated Vulnerability Explanation Annotation: Traditional open-source vulnerability datasets mainly contain source code, vulnerability labels, CWE types, and commit messages.
+However, they often lack detailed explanations of the vulnerability logic within the source code, posing a significant challenge for vulnerability detection techniques that aim to provide meaningful explanations for detected issues.
+Manually annotating real-world vulnerable code explanations requires extensive software development experience and a deep understanding of software vulnerabilities, which incurs high labor and time costs [36].
+As the scale of manually annotated data remains a significant challenge, an increasing number of researchers are leveraging the powerful generative capabilities of LLMs to synthesize data.
+This approach has been validated in various domains through improvements in downstream task performance metrics.
+However, in the specialized field of vulnerability detection, where domain expertise is critical, the feasibility of synthesizing data has yet to be fully validated.
+To address this challenge, we introduce an innovative automated vulnerability explanation annotation method based on prompt engineering with LLMs.
+This method capitalizes on the contextual learning and instruction-following capabilities of LLMs, enabling large-scale, high-quality automated synthesis of vulnerability explanations.
+Our approach decomposes the explanation task into three sub-goals: (1) vulnerability discrimination, (2) identifying the location of the vulnerability in the code, and (3) providing a specific explanation.
+The model is guided by instruction-based prompt templates, paired with well-annotated examples, which stimulate its contextual learning capabilities, ensuring that the generated explanations are both accurate and informative.
+We implement this annotation process using GPT-3.5 [37], accessed via the API provided by OpenAI [38] due to its balance of annotation cost and efficiency.
+As part of our experimental study, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively.
+This effort fills a significant gap in the availability of vulnerability explanation annotations, enabling more effective training and evaluation of vulnerability detection and explanation models.
+To validate the effectiveness of the synthesized data, we indirectly evaluated the performance of the fine-tuned model in vulnerability detection tasks across multiple scenarios.
+Additionally, we propose evaluation metrics for vulnerability explanation quality, along with both manual and automated evaluation methods, to directly assess the quality of the generated explanations.
+❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated annotation of vulnerability explanations for open-source data creates a large-scale dataset, effectively addressing data bottlenecks in the fine-tuning process for vulnerability detection and explanation.
+To enhance the detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune general LLMs to specialize in detecting and explaining specific types of vulnerabilities in real-world code.
+Instruction-based prompts are used to guide these tasks, helping LLMs to accurately understand the objectives and generate standardized outputs.
+To minimize the computational cost of the fine-tuning process, we employ the parameter-efficient fine-tuning method LoRA [12], which significantly reduces both time and space requirements.
+❹ Evaluation of Generative Vulnerability Explanation: The limited research on model-generated vulnerability explanations underscores the need for robust evaluation methods tailored to LLM-generated outputs.
+Similar to annotation challenges, manual evaluation requires considerable human and time resources.
+Moreover, effective vulnerability explanations must be assessed across multiple dimensions to ensure they offer practical value to developers in real-world scenarios.
+To address this, we propose an evaluation framework based on three critical dimensions: accuracy, clarity, and actionability.
+These dimensions collectively measure the correctness, comprehensibility, and practical applicability of the generated explanations.
+To enhance evaluation efficiency, we introduce an automated evaluation method based on LLMs.
+This method leverages prompt engineering to enhance evaluation efficiency while ensuring reliable assessments.
+Additionally, expert manual verification is employed to validate both the quality of the outputs from the specialized vulnerability explanation model and the feasibility of the LLM-based automated evaluation scheme.
 III-A Vulnerability Interpretation Enhancement Prompting
-Despite the robust code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning tasks that require deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities.
-These challenges manifest as insufficient detection accuracy and vague vulnerability analysis.
-To guide LLMs toward better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs.
-By leveraging prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
-Figure 3 illustrates the design of our prompt templates.
-These templates consist of four main components: task description, specific instructions, generation examples, and sample inputs.
-(1) Task description: Provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format.
-This helps the LLM understand task requirements and grasp the background knowledge of the vulnerabilities.
-(2) Specific instructions: Include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output.
-These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis.
-(3) Generation examples: Provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals.
-(4) Sample input: Used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
-Figure 3: Prompting template of Automated Vulnerability Interpretation Labelling.
-Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, resulting in well-performing fine-tuned specialized vulnerability detection and explanation models.
+Despite the robust code understanding and analysis capabilities of LLMs, they face significant challenges in complex reasoning tasks that demand a deep understanding of code, advanced reasoning abilities, and specialized knowledge of vulnerabilities.
+These challenges often result in insufficient detection accuracy and vague vulnerability analyses.
+To address these limitations, we integrate instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs.
+By leveraging prior knowledge from open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively enabling the application of LLMs in vulnerability detection and explanation tasks across all critical stages of the framework.
+Figure 3 illustrates the structure of our prompt templates.
+These templates are composed of four main components: task description, specific instructions, generation examples, and sample input.
+(1) Task description: Specifies the template for the current vulnerability detection and explanation task, including details about the types of vulnerabilities being addressed and the basic input-output format.
+This component provides the LLM with a clear understanding of task requirements and relevant background knowledge.
+(2) Specific instructions: Define the required input-output format for the LLM, such as step-by-step output guidelines, the range of vulnerabilities to focus on, and output length constraints.
+These instructions utilize the LLM’s instruction-following ability to ensure standardized and uniform outputs, which facilitates subsequent processing and analysis.
+(3) Generation examples: Contain manually curated samples of vulnerable code snippets paired with corresponding explanation data.
+These examples help the LLM better comprehend the task’s goals and improve the quality of generated outputs.
+(4) Sample input: Includes the code to be analyzed and, during the annotation phase, incorporates associated labels and supplementary information such as CWE types, CVE descriptions, or commit messages.
+Figure 3: Prompt template of Explanation Annotation.
+Through rigorous testing and evaluation across various vulnerability detection and explanation tasks, we validate that these prompt templates effectively achieve their intended objectives at all stages.
+The resulting fine-tuned models demonstrate improved performance in specialized vulnerability detection and explanation tasks.
 III-B Key Code Extraction Based on Chain-of-Thought (CoT)
-One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets.
-Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs.
-However, these methods are often difficult to apply in diverse vulnerability detection scenarios.
-To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction.
-This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code.
-The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities through a thought chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
-The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on.
-After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT.
-This approach guides the model step by step to reference the key statements to complete the vulnerability detection, localization, and explanation.
+One of the key challenges in explaining code vulnerabilities lies in accurately identifying and extracting critical code segments—referred to as key code—from lengthy code snippets.
+Key code typically includes statements, variable structures, and other components that are closely related to the vulnerability type under investigation.
+These elements often represent the root cause or propagation paths of vulnerabilities and provide essential context for understanding the issue.
+However, traditional static analysis techniques, such as abstract syntax trees, program dependence graphs, and control flow graphs, often fail to generalize effectively in diverse and complex vulnerability detection scenarios.
+To address these challenges and improve the capability of large models to analyze complex code structures, we propose a novel Chain-of-Thought (CoT) enhancement method that incorporates automated key code extraction.
+This approach enables the model to focus on the most relevant parts of the code, thereby improving accuracy and interpretability.
+Using the prompt template illustrated in Figure LABEL:fig:prompt_key, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type.
+These extracted statements form a structured CoT, guiding the fine-tuned model through a step-by-step process for targeted vulnerability detection, location, and explanation.
+For example, in a buffer overflow scenario, key code may include array declarations, bounds-checking statements, or pointer dereferences that contribute directly to the vulnerability.
+By integrating these extracted key statements into the prompts during fine-tuning, the model can systematically reference this crucial information, enhancing both the interpretability and precision of the vulnerability explanation process.
+This CoT-based key code extraction methodology ensures that the model is not only detecting vulnerabilities more accurately but also providing detailed, context-rich explanations that developers can use to understand and address these issues effectively.
+Figure 4: Prompt template of Key Code Extraction.
 IV Results
@@
 In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
-IV-A1 Experimental Setup
+IV-A1 Experimental Setup.
 Dataset.
@@
 Backbone LLMs.
-We use the Codellama [39] and Llama3 as our backbone LLMs for fine-tuning.
+In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [39] as our primary model for the experiments discussed in this chapter.
 CodeLlama is initialized with the weights of Llama2 [8] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities.
-The Llama-family LLMs have been widely used by prior studies related to LLMs for software engineering [41, 42, 43].
-In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [44] as our primary model for the experiments discussed in this chapter.
-For comparison purposes, we also considered the 7B version and the newly released Llama3-8B-Instruct in the discussion section.
-Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design four sub-research questions:
+Experimental Setting.
+To comprehensively evaluate the fine-tuned vulnerability LLMs, we design three sub-research questions:
   • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
-  • (RQ1.2) What impact does the inclusion of other types of vulnerability data in training have on the detection performance of the target type of vulnerability?
-  • (RQ1.3) How effective is a unified model in detecting multiple types of vulnerabilities?
-  • (RQ1.4) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
+  • (RQ1.2) How effective is the model in detecting multiple types of vulnerabilities simultaneously?
+  • (RQ1.3) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
 To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics.
@@
 For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1.
-To test the detection accuracy, we selected CodeT5 [45] and CodeBERT [46] as baselines, which have been commonly used by prior studies in software engineering [47, 48].
-In particular, we fine-tune them for classification tasks by adding a linear classification layer.
+To test detection accuracy, we selected CodeT5 [40] and CodeBERT [41] as baselines and fine-tuned them for classification tasks by adding a linear classification layer.
+These models are chosen for two primary reasons: (1) As pre-trained code models, CodeT5 and CodeBERT have demonstrated consistent and superior performance on prior vulnerability detection benchmarks [16, 14], making them reliable baselines.
+(2) Their inclusion allows us to assess whether fine-tuned LLMs offer an advantage over existing pre-trained code models in terms of detection accuracy.
+It is important to note that the focus of this study is not on developing the most optimal detection model but rather on evaluating the explanation capabilities of fine-tuned LLMs.
+Therefore, using these two baselines provides a meaningful comparison without detracting from the primary objective of assessing the interpretability of LLM-generated explanations.
 Implementation Details.
-We implement our approach using the Transformers [17] and PEFT [49] libraries with the PyTorch platform.
-All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs platforms, with the token length limit set to 2048.
-We use the AdamW optimizer and train the models for 3 epochs.
-For LoRA configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
-IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model
-To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type.
-The sample size of each type is shown in Table II.
-In the training and inference prompts, we specify the type of vulnerability to help narrow the scope of detection.
-During training, we use explanations annotated by GPT-3.5, which include vulnerability location information and specific details, as the target output for the model to perform detection and explanation tasks.
-The inputs for the model during both training and inference are code snippets from the given examples.
-As a generative model, we conduct a binary classification task based on the semantics of the generated text.
-For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it outputs a predefined fixed pattern (e.g., “There are no security issues”).
+We implement our approach using the Transformers[13] and PEFT[42] libraries on the PyTorch platform.
+All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs, with a token length limit set to 2048.
+The models are trained using the AdamW optimizer for 3 epochs with a batch size of 2 and data precision set to float16 to optimize memory usage and computational efficiency.
+We set the learning rate to 0.0003 and weight decay to 0.01.
+For the LoRA configuration, we set target_modules to {q_proj, k_proj, v_proj, o_proj} for SeVC and {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj} for DiverseVul.
+The LoRA rank is set to 16, the LoRA scaling factor to 16, and the dropout rate to 0.05.
+For fine-tuning CodeBERT and CodeT5, we use the AdamW optimizer with the learning rate of 1e-5 and the weight decay of 1e-2, and train the models for 3 epochs.
+All other parameters are set to their default values.
+IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model.
+To evaluate the detection capability for a single type of vulnerability, we fine-tune a separate model for each vulnerability type on SeVC, resulting in specialized models tailored to individual vulnerability types.
+The sample sizes of each type is shown in Table II.
+In both training and inference prompts, explicitly specify the target vulnerability type to narrow the scope of detection and improve precision.
+During training, the model utilizes explanations annotated by GPT-3.5, which include detailed information about the vulnerability location and its specific characteristics, as the target output for performing both detection and explanation tasks.
+The model’s input during both training and inference consists of code snippets from the corresponding dataset examples.
+As a generative model, the task is formulated as a binary classification based on the semantics of the generated text.
+For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it produces a predefined fixed pattern (e.g., “There are no security issues”).
 TABLE II: Performances Comparison of fine-tuning on SeVC(RQ1.1).
@@
 Experimental Results.
-We first compared the impact of fine-tuning on the detection accuracy of CodeLlama.
-As shown in Table II, CodeLlama-13B-Instruct lacks precise vulnerability identification capabilities.
-Fine-tuning significantly improves the detection accuracy.
-Therefore, in subsequent tasks, we no longer compare with the original CodeLlama.
-From Table III we can observe that our generative model still achieves comparable performance to other fine-tuned classification models.
-On the one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type.
-On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs.
-If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
-IV-A3 RQ1.2: Detection with specified vulnerability type using an all-in-one model
-To explore the impact of training with data from different types of vulnerabilities on the detection effectiveness for target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning a model on four types of vulnerabilities and then performing individual detection for each type.
-Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection.
-During inference, we also provide the vulnerability type information for each example.
-The difference between RQ1.1 and RQ1.2 is that in RQ1.1, we fine-tune a dedicated model for each type of vulnerability, whereas in RQ1.2, we fine-tune an all-in-one model and use prompts to distinguish the type of vulnerability of interest.
-TABLE IV: Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
-Accuracy Precision Recall F1
-Type Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one
-API 93.0% 97.1% 91.6% 97.4% 94.5% 96.8% 93.0% 97.1%
-Arithmetic 91.7% 96.6% 90.3% 97.6% 93.6% 95.6% 91.9% 96.6%
-Pointer 92.7% 97.1% 93.7% 97.6% 91.2% 96.6% 92.4% 97.1%
-Array 92.7% 96.6% 95.3% 97.0% 89.7% 96.3% 92.4% 96.6%
-Average 92.5% 96.9% 92.7% 97.4% 92.2% 96.3% 92.4% 96.9%
-Experimental Results.
-We present the results of the fine-tuned all-in-one model in Table IV.
-In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities.
-This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities.
-This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs.
-Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
-IV-A4 RQ1.3: Detection with an identification of the vulnerability type
-In real software development environments, the vulnerability risks faced are often diverse, making it difficult to preemptively sense the potential types of vulnerabilities in the samples to be detected.
-This requires the large model to accurately identify and classify multiple types of vulnerabilities and provide specific explanations.
-To test the accuracy of a unified model in this context, we conduct a multi-class vulnerability detection fine-tuning on SeVC with all samples.
-Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including non-vulnerable and the four types of vulnerabilities in SeVC.
-In the training and inference prompts, we no longer specify the type of vulnerability.
-Instead, we adopt a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
-TABLE V: Performances of Multi-Type Detection on SeVC(RQ1.3).
+We first evaluate the impact of fine-tuning on the detection accuracy of CodeLlama.
+As shown in Table II, the original CodeLlama-13B-Instruct model lacks precise vulnerability identification capabilities, demonstrating suboptimal performance in detecting specific types of vulnerabilities.
+However, fine-tuning significantly improves detection accuracy.
+Consequently, we exclude the original CodeLlama from subsequent comparisons.
+From Table III, we observe that our generative model achieves performance comparable to other fine-tuned classification models.
+This result highlights two key findings: (1) Our approach effectively enhances the model’s understanding of the specific vulnerability type, enabling it to capture critical patterns for accurate detection.
+(2) The detection task for a single type of vulnerability is relatively less challenging for LLMs.
+In practical scenarios where the focus is on a limited set of vulnerability types, fine-tuning with target-specific data can yield superior performance.
+IV-A3 RQ1.2: Detection with identification of multiple vulnerability types.
+In real software development environments, vulnerability risks are often diverse, posing challenges in accurately predicting the specific types of vulnerabilities present in the code.
+This necessitates a model capable of simultaneously detecting, classifying, and explaining various vulnerability types.
+To evaluate the effectiveness of such a model, we train a single model on SeVC for multi-class vulnerability detection using all available samples.
+Specifically, a ‘[type]’ tag is appended to the model’s output to classify the detected vulnerabilities, encompassing non-vulnerable code and the four distinct types of vulnerabilities present in SeVC.
+Unlike the previous section, where each model is trained to detect a single type of vulnerability, here we use a unified approach to enable one model to handle all types concurrently.
+During training and inference, the specific vulnerability type of detected code will not appear in the prompt.
+Instead, a task description encompassing all four vulnerability types in the current multi-class scenario is provided, enabling the model to develop the capability to distinguish between multiple types of vulnerabilities.
+TABLE IV: Performances of Multi-Type Detection on SeVC (RQ1.2).
 Precision Recall F1
@@
 Experimental Results.
-We present the metrics for each type for the multi-class vulnerability detection task in Table V.
-Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code.
-(2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario.
-(3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types.
-The other three types exhibit a significant degree of confusion and are difficult to identify.
-This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient.
-The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
-IV-A5 RQ1.4: Identification of the vulnerability type in datasets with more vulnerability types
-In real development environments, security risks stem from a wide variety of project types and complex code structures, often involving a broader range of vulnerability models.
-To better resemble real development conditions, we conduct a multi-label classification task on the top 10 types of CWE in the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels).
-Specifically, we modify the task description to focus on CWE-type vulnerability detection and explanation tasks, adding a ‘[CWE]’ tag to generate a list of CWEs related to the target code.
-Additionally, we incorporate CWE descriptions as part of the output explanation to help the model better understand the meaning of each CWE type and utilize these descriptions for vulnerability analysis and identification.
-TABLE VI: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.4).
-Metric CWE ID Overall (Micro)
+The metrics for each type in the multi-class vulnerability detection task are presented in Table IV.
+Based on the experimental results, we observe the following: (1)The non-vulnerable code category exhibits high precision and recall, indicating that the model maintains strong capability in identifying non-vulnerable code even in a multi-type scenario.
+(2) Compared to detection results where the type of interest is provided, the overall performance shows varying degrees of decline, indicating that the model faces greater difficulty in distinguishing between vulnerability types.
+(3) Among the vulnerability types, Arithmetic Expression achieves relatively better performance, whereas the remaining three types demonstrate significant confusion, making them harder to identify.
+These findings indicate that, while the model is effective at recognizing non-vulnerable code, its ability to differentiate between multiple vulnerability types is limited, particularly when certain types are underrepresented in the training data.
+This highlights the critical role of balanced and representative training datasets in improving the model’s performance across all vulnerability types.
+IV-A4 RQ1.3: Identification of a greater variety of vulnerability types.
+In real-world software development environments, security risks arise from a diverse range of project types and complex code structures, often involving a broader spectrum of vulnerability categories.
+To better reflect these real-world conditions, we conduct a multi-label classification task using the top 10 CWE categories in the real-world vulnerability dataset DiverseVul, where a subset of the code samples is annotated with multiple CWE labels.
+Specifically, the task description is modified to emphasize CWE-type vulnerability detection and explanation, with the addition of a ‘[CWE]’ tag to generate a list of CWE categories relevant to the target code.
+Furthermore, CWE descriptions are incorporated as part of the output explanation, aiding the model in comprehending the context and semantics of each CWE type and leveraging these descriptions for vulnerability analysis and identification.
+TABLE V: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.3).
+Metric CWE ID Overall (Micro-F1)
 787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
@@
 Experimental Results.
-The experimental results are presented in Table VI.
-From the results, we find that after fine-tuning, the specialized vulnerability models achieve good detection accuracy for the 10 types of CWE, demonstrating their ability to grasp the pattern information for each CWE type.
-Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories.
-Each vulnerability type has a more fine-grained and explicit definition, avoiding potential category association issues in SeVC and resulting in better performance.
-This validates the feasibility of using LLMs for vulnerability detection and explanation in real software development environments, which is illustrated in the following explanation example.
+The experimental results are presented in Table V.
+The findings demonstrate that after fine-tuning, the specialized vulnerability models achieved high detection accuracy across the 10 CWE categories, highlighting their capability to capture the distinctive patterns associated with each CWE type.
+Unlike the multi-class task in SeVC, the multi-label classification in DiverseVul does not require the model to distinguish between mutually exclusive categories.
+Instead, each CWE type has a more granular and explicit definition, mitigating the category association challenges observed in SeVC and thereby yielding improved performance.
+These results validate the practicality of employing LLMs for vulnerability detection and explanation in real-world software development scenarios.
+The following example of a generated explanation further illustrates the model’s capacity to analyze and identify CWE-based vulnerabilities effectively.
 1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.
@@
 The vulnerability is related to CWE-476: NULL Pointer Dereference.
-TABLE VII: Example of Generated Explanation on DiverseVul.
+TABLE VI: Example of Generated Explanation on DiverseVul.
 Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
 IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
-In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities.
-Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanations.
-The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content.
-Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
-Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics: Accuracy, Clarity, and Actionability, as the evaluation criteria for vulnerability explanations.
-We use both manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
-IV-B1 Experimental Setup
-In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT-3.5.
-Due to the large scale of the dataset we use and the high evaluation cost, we randomly sample the data based on a 95% confidence level and a 5% confidence interval [50] following prior studies [51, 52, 53] for the fine-tuned model’s correctly classified instances.
-We obtain 384 and 224 samples from SeVC and DiverseVul, respectively.
-This sampling method ensures the statistical validity and representativeness of our evaluation.
-Our evaluation includes two parts: (1) Manual review: Two authors of this paper, both of whom have extensive research experience in software vulnerabilities and software development, independently conduct manual checks on the samples, scoring them according to the criteria.
-In cases of disagreement, discussions are held until a consensus is reached.
-The value of Cohen’s Kappa [54] in this process is 0.76, indicating substantial agreement.
-This rigorous manual review process ensures the reliability and accuracy of our evaluation.
-(2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion.
-We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction.
-This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.
-Evaluation Metrics. The specific meanings and requirements of the three metrics are as follows:
-  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring that the provided details are factually correct and relevant to the detected vulnerability.
-  • Clarity: The information should be presented in a clear, understandable manner, and structured in a way that facilitates easy comprehension by software developers.
-  • Actionability: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
-IV-B2 Experimental Results
-We present the results of our dual evaluation process in Table VIII, the suffix Gen.
-refers to the explanation generated by LLMVulExp and Ann.
-refers to the explanation annotated by GPT-3.5.
+In this RQ, we explore the proficiency of the LLMVulExp in explaining the detected vulnerabilities.
+While numerous evaluation metrics exist for assessing text generation by LLMs [43, 44], there is currently no comprehensive framework tailored to evaluating vulnerability explanations.
+The explanatory content generated by the model should assist software developers in identifying and mitigating potential vulnerabilities by ensuring the accuracy of the analysis, maintaining readability and conciseness, and providing actionable suggestions for remediation.
+The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a ”correct” or ”complete” explanation.
+This inherent subjectivity poses significant challenges for effective evaluation.
+Therefore, the evaluation of explanations must be contextualized within the specific goals of the vulnerability explanation task.
+By summarizing and generalizing the “helpfulness” of the generated text, we assess explanation quality from three critical perspectives: Accuracy, Clarity, and Actionability.
+IV-B1 Experimental Setup.
+In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the SeVC and DiverseVul test sets, alongside ground truth annotations generated using GPT-3.5.
+Given the large scale of the dataset and the high cost of evaluation, we employ a random sampling approach based on a 95% confidence level and a 5% confidence interval [45].
+This yields 384 and 224 samples from SeVC and DiverseVul, respectively, ensuring statistical validity and representativeness in our evaluation.
+Our evaluation includes two parts: (1) Manual review: Two authors, with extensive research experience in software vulnerabilities and development, independently review the sampled data, scoring them based on predefined criteria.
+Disagreements are resolved through discussion until a consensus is reached.
+The Cohen’s Kappa value [46] for inter-reviewer agreement is 0.76, indicating substantial consistency and reliability.
+(2) Automated review using LLM: GPT-3.5 is used for automated evaluation.
+A detailed prompt is provided, describing the vulnerability explanation evaluation task and the specific criteria.
+The LLM is instructed to output scores for the three evaluation metrics, with 1 indicating satisfaction and 0 indicating non-satisfaction.
+This dual evaluation approach provides a robust assessment of the quality of the explanations generated by the fine-tuned model and demonstrates the feasibility of using LLMs for scalable evaluation tasks in the future.
+Evaluation Metrics. The three evaluation metrics used in this study are defined as follows:
+  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring the details are factually correct and relevant to the detected vulnerability.
+  • Clarity: The explanation should be clear, concise, and structured in a way that facilitates easy comprehension by software developers.
+  • Actionability: The explanation include actionable suggestions for remediation, offering practical guidance for addressing the identified vulnerability.
+IV-B2 Experimental Results.
+We present the results of our dual evaluation process in Table VII.
+The suffix Gen.
+refers to explanations generated by LLMVulExp, while Ann.
+refers to annotations from GPT-3.5.
 All-Pos.
-refers to the results that meet the requirements of all of the three metrics.
-Combining the results from manual and LLM evaluations, we have made the following observations:
-(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen.
-for both the manual review and automated review) in the explanation.
-They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
-(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively.
-The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
-(3) Actionability of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
-(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann.
-and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data.
-It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
-(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review.
-For SeVC and DiverseVul, the results of All-Pos.
-comparing manual review and automated review are 76.0% vs.
-74.0% and 80.1% vs.
-71.9%, respectively.
-This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text.
-In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
-These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation.
-The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations.
-The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities.
-The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis.
-Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
-TABLE VIII: Distribution (%) of Vulnerability Explanation Review Results (RQ2).
+represents cases meeting all three evaluation metrics.
+Key findings include:
+(1) Accuracy of Explanations: LLMVulExp achieves high accuracy (e.g., over 90.0% for SeVC-Gen.
+in both manual and automated reviews), pinpointing vulnerability risks and code locations.
+This underscores the importance of accurate vulnerability type identification.
+(2) Clarity of Explanations: The clarity scores are 81.4% for SeVC and 94.1% for DiverseVul, indicating the model effectively reduces cognitive barriers for developers analyzing vulnerabilities.
+(3) Actionability of Explanations: Explanations from LLMVulExp include actionable suggestions for code modifications, achieving 93.4% for SeVC and 80.5% for DiverseVul.
+This highlights the model’s potential to support practical remediation efforts.
+(4) Effectiveness of Annotation Method: Results from SeVC-Ann.
+and DIV-Ann.
+validate the effectiveness of our annotation method in generating high-quality explanatory data, reducing subsequent annotation costs.
+(5) Potential of LLM in Automated Assessment: LLM-based automated evaluations closely align with manual reviews.
+For All-Pos., manual vs.
+automated results are 76.0% vs.
+74.0% for SeVC and 80.1% vs.
+71.9% for DiverseVul.
+This demonstrates LLMs’ potential to scale evaluation tasks efficiently.
+While the proposed approach demonstrates strong performance, it has inherent limitations.
+Fully correct explanations cannot be guaranteed due to the complexity of vulnerability analysis, and manual evaluations are subject to reviewer bias.
+Despite these challenges, the dual-validation strategy combining manual and automated reviews offers a practical and effective solution for evaluating vulnerability explanations, balancing accuracy with scalability and cost-effectiveness.
+TABLE VII: Proportion (%) of Vulnerability Explanation Review Results (RQ2).
 Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
@@
 Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
-TABLE IX: Example of Generated Explanation on SeVC.
-IV-B3 Case Study
-To more intuitively illustrate the importance of evaluating the quality of vulnerability explanations, we analyze an example generated by our model.
-As shown in Figure IV-B2, this vulnerability code is sourced from SeVC and pertains to pointer-type vulnerabilities, demonstrating incorrect memory deallocation in C language using the delete operator.
-The explanation adheres to our output paradigm, generating vulnerability detection, location, and detailed information.
-The provided explanation correctly identifies the core issue: the data pointer is allocated on the stack using the ALLOCA function but is improperly freed using the delete operator.
-The explanation points out that the delete operator should only be used to free memory allocated on the heap via the new operator.
-Using it to free stack-allocated memory can lead to undefined behavior and potential crashes, meeting the accuracy requirement.
-Additionally, the explanation offers recommendations for addressing this vulnerability, such as using the free function to release stack-allocated memory or using the new operator for heap allocation, and ensuring pointer safety.
-This satisfies the actionability requirement.
-The explanation is clear, comprehensive, logically coherent, and easy to read, meeting the clarity requirement.
-During the review, we also observe some instances with insufficient quality of explanations.
-Issues include failure to identify the main vulnerability, vague location, verbose expression, and lack of remediation suggestions.
-Hallucinations in explanation generation are relatively rare but primarily involved memorized descriptions of similar code samples from the training set and location confusion.
-Overall, the fine-tuned model demonstrates a certain capability for generating vulnerability explanations.
-Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Actionability.
-We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
+TABLE VIII: Example of Generated Explanation on SeVC.
+IV-B3 Illustrative Example.
+To highlight the importance of evaluating vulnerability explanations, we analyze an example from SeVC.
+The sample pertains to a pointer-type vulnerability caused by incorrect memory deallocation in C code.
+As shown in Figure IV-B2, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function.
+The explanation provides actionable suggestions, such as ensuring pointer safety by using the appropriate deallocation methods.
+It is clear, logically structured, and actionable, satisfying all three evaluation criteria.
+Despite these strengths, some generated explanations exhibit issues, such as vague locations, verbose descriptions, or lack of actionable suggestions.
+While rare, hallucinations may occur, typically involving memorized examples or misinterpreted contexts.
+Overall, the fine-tuned model demonstrates robust capabilities for generating high-quality vulnerability explanations.
+Summary of RQ2: We find that LLMVulExp is capable of generating vulnerability explanations with high levels of Accuracy, Clarity, and Actionability.
+Additionally, we explore the potential of leveraging LLMs for automated data annotation, significantly reducing the reliance on manual effort.
 IV-C RQ3: How do explanations affect the results of vulnerability detection?
-Considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1.
-This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs compared to fine-tuning specialized detection models?
-In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory information.
+Building on the detection capabilities evaluated in RQ1, we consider the detection task as a foundational step in the broader explanation task.
+This raises an important question: How does fine-tuning for vulnerability explanation influence the detection performance of LLMs compared to models fine-tuned solely for detection?
+In this RQ, we experimentally examine the impact of incorporating explanatory information on vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory data.
 Experimental Setup.
-We conduct our research as ablation studies under the settings of multi-type vulnerability detection scenarios in SeVC (RQ1.3) and multi-label CWE detection scenarios in DiverseVul (RQ1.4).
-We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt.
-By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model.
-We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
-Figure 4: Performance Comparison between Explainer and Detector.
+We conduct ablation studies under the multi-type vulnerability detection scenario in SeVC (RQ1.2) and the multi-label CWE detection scenario in DiverseVul (RQ1.3).
+To isolate the effect of explanatory information, we transform the vulnerability explanation fine-tuning task into a detection-only task by removing explanatory data from annotations and adjusting task descriptions and prompts.
+The resulting LLMs output only vulnerability types, creating a detection-focused model (referred to as Detector).
+This is compared against the performance of the original explanation-focused model (referred to as Explainer).
+Figure 5: Performance of Explainer and Detector.
 Experimental Results.
-We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 4.
-The prefix of “Mi” and “Ma” refers to Micro and Macro, respectively.
-Based on the experimental results, incorporating vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance.
-For example, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% for both on DiverseVul.
-In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
-These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities.
-The model’s detection capability remains intact with an additional focus on explanation capabilities.
-Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns.
-This suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance.
-Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
-Summary of RQ3: Integrating vulnerability explanation into the fine-tuning process does not compromise detection capabilities and may even lead to improved performance for certain vulnerability scenarios.
+The comparative results between Explainer and Detector are illustrated in Figure 5, where ”Mi” and ”Ma” denote Micro and Macro metrics, respectively.
+Incorporating explanatory data into the fine-tuning process did not significantly reduce detection performance.
+For instance, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while both achieve 74.8% on DiverseVul.
+Interestingly, certain metrics improved when explanatory data was included, such as the Macro-F1 score on DiverseVul, where the Explainer outperformed the Detector by nearly 2% (73.2% vs.
+71.4%).
+These findings demonstrate that the vulnerability explanation task can coexist with the detection task without compromising the model’s detection accuracy.
+The Explainer retains strong detection performance while gaining additional explanatory capabilities.
+Furthermore, automatically annotated explanatory data introduces domain-specific knowledge, enhancing the model’s ability to understand and identify diverse vulnerability patterns.
+In summary, enhancing a model’s explanatory capabilities does not come at the cost of detection performance.
+In some cases, well-annotated explanatory data can even improve the model’s overall understanding and effectiveness across both tasks.
+Summary of RQ3: Integrating vulnerability explanations into the fine-tuning process preserves detection capabilities and may even enhance performance in specific vulnerability scenarios.
 IV-D RQ4: How does the key code extraction impact detection performance?
-In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to vulnerability and further examine the code with a specific focus.
-To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
+In this RQ, we investigate whether LLMs can identify key code segments prone to vulnerabilities and enhance detection by focusing on these critical sections.
+To explore this, we annotate key code segments in the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the annotated key code information.
 Experimental Setup.
-Similar to the process of annotation discussed in Section III, we use GPT-3.5 to extract the key code.
-To prevent label leakage, the code vulnerability type tags were not visible during the annotation process.
-For training data that does not output the original code statements, we will nullify its key code information.
-If it is test data, it will not be included in the evaluation.
-During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
-TABLE X: F1 Scores of Key Code Extraction on SeVC (RQ4).
+As outlined in Section III, we employ GPT-3.5 to annotate key code segments closely related to vulnerabilities.
+To avoid label leakage, vulnerability-type tags are concealed during the annotation process.
+For training samples where key code information was unavailable, the annotations are nullified.
+Similarly, test samples without key code annotations are excluded from the evaluation.
+During fine-tuning and inference, the prompts are modified to emphasize task descriptions and instructions, guiding the model to focus on the extracted key code for vulnerability detection.
+TABLE IX: F1 of Key Code Extract. on SeVC: (RQ4).
 Type Single-Type(RQ1.1) Multi-Type(RQ1.3)
@@
 Experimental Results.
-We present the experimental results of utilizing Key Code Extraction on SeVC in Table X and DiverseVul in Figure 5.
+We present the experimental results of utilizing Key Code Extraction on SeVC in Table IX and DiverseVul in Figure 6.
 We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type).
@@
 This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models.
-On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results.
-On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio.
-This makes it easier for the model to learn and detect patterns associated with vulnerabilities.
-This method does not rely on manual feature extraction, making it broadly applicable.
-Our experimental results preliminarily demonstrate the feasibility of key code extraction through LLMs.
-Figure 5: Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent the CWE IDs.
-Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
+Key code extraction offers two major benefits: (1) Enhanced Focus: By isolating the most relevant code segments, the model can focus on critical information, leading to more accurate detection results.
+(2) Noise Reduction: Eliminating irrelevant code improves the signal-to-noise ratio, enabling the model to better learn and detect vulnerability patterns.
+Moreover, this approach relies on automated annotation rather than manual feature extraction, making it broadly applicable and scalable.
+Our experimental results preliminarily validate the feasibility and utility of key code extraction through LLMs.
+Figure 6: Performance of Key Code on DiverseVul(RQ4).
+Summary of RQ4: By guiding LLMs to focus on key code, the performance of vulnerability detection can be significantly improved(e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
 V Discussion
@@
 Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
-Figure 6: Performance of Multi-Type vulnerability detection using LLMs with different sizes.
+Figure 7: Performance of different base LLMs.
 Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks.
@@
 Construct Validity.
-We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Actionability.
-While these metrics effectively characterize important aspects of vulnerability explanations, they may not comprehensively cover all features.
-More comprehensive and refined evaluation metrics for vulnerability explanations require further research.
-To mitigate potential bias, we employed both manual and automated reviews using LLMs.
-Two authors independently annotated the explanations and resolved any discrepancies until a consensus was reached.
-The value of Cohen’s Kappa [54] in this process was 0.76, indicating substantial agreement.
-This rigorous approach helps ensure the reliability and validity of our evaluation.
+We evaluate explanatory capability using three effective metrics, but these metrics may not cover all aspects of explanations.
+To mitigate bias, we use both manual and automated reviews with LLMs.
+Two authors independently annotate the explanations, resolving discrepancies until consensus is reached, achieving a Cohen’s Kappa [46] value of 0.76, indicating substantial agreement.
 Internal Validity.
-Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states.
-The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results.
-However, we conducted multiple trials to validate the stability of our conclusions for each experimental setup.
-Although the number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited, and due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version, we ensured that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
+Due to the high computational cost of fine-tuning and inference evaluation, we couldn’t conduct experiments under different data splits and randomization states.
+Although we limit the number of hyperparameter trials for LoRa configuration and didn’t fine-tune larger models like the 34B version due to GPU constraints, these limitations don’t undermine our objective: exploring the potential of LLMs for vulnerability detection and explanation.
 External Validity.
-We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research.
-We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks.
-However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models.
-Additionally, we conducted experiments using two differently configured open-source project-based vulnerability datasets.
-Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments.
-Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments.
-Nonetheless, our careful selection and thorough evaluation process aimed to ensure that our findings remain relevant and insightful within the scope of our study.
+We choose the advanced open-source code LLM CodeLlama and evaluate different versions as well as the recently released Llama3 on two distinct datasets.
+These datasets, while containing a significant amount of vulnerable code, include only a few CWE types and mostly generic vulnerabilities, which do not fully reflect the diversity and complexity of vulnerabilities found in real-world development environments.
+Although these datasets contain substantial vulnerable code, they differ from real development environments, posing a risk of limited generalization.
 VII Conclusion

### v4

@@
-mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt
+Towards Explainable Vulnerability Detection with Large Language Models
 Abstract
-Software vulnerabilities pose significant risks to the security and integrity of software systems.
-Although prior studies have explored vulnerability detection using deep learning and pre-trained models, these approaches often fail to provide the detailed explanations necessary for developers to understand and remediate vulnerabilities effectively.
-The advent of large language models (LLMs) has introduced transformative potential due to their advanced generative capabilities and ability to comprehend complex contexts, offering new possibilities for addressing these challenges.
-In this paper, we propose LLMVulExp, an automated framework designed to specialize LLMs for the dual tasks of vulnerability detection and explanation.
-To address the challenges of acquiring high-quality annotated data and injecting domain-specific knowledge, LLMVulExp leverages prompt-based techniques for annotating vulnerability explanations and fine-tunes LLMs using instruction tuning with Low-Rank Adaptation (LoRA), enabling LLMVulExp to detect vulnerability types in code while generating detailed explanations, including the cause, location, and repair suggestions.
-Additionally, we employ a Chain-of-Thought (CoT) based key code extraction strategy to focus LLMs on analyzing vulnerability-prone code, further enhancing detection accuracy and explanatory depth.
-Our experimental results demonstrate that LLMVulExp achieves over a 90% F1 score on the SeVC dataset, effectively combining high detection accuracy with actionable and coherent explanations.
-This study highlights the feasibility of utilizing LLMs for real-world vulnerability detection and explanation tasks, providing critical insights into their adaptation and application in software security.
-11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
+Automated vulnerability detection tools often lack the explainability needed to assist developer remediation.
+While Large Language Models (LLMs) are promising, creating explainable detectors faces two key challenges: acquiring high-quality explanation data at scale and instilling models with autonomous, end-to-end reasoning capabilities.
+To address these challenges, we propose LLMVulExp, an innovative “teacher-student” framework.
+Our core contribution is a pipeline that resolves the data bottleneck: we guide a powerful “teacher” LLM to generate a large-scale corpus of high-quality explanations for known vulnerabilities.
+This data is then used to efficiently train a “student” LLM, via instruction tuning and LoRA, to autonomously detect and explain vulnerabilities in unseen code.
+Our second key contribution is a novel evaluation protocol for explanations, featuring metrics for Accuracy, Clarity, and Actionability, alongside an automated “LLM-as-a-judge” method.
+This work thus presents a complete and validated end-to-end framework for building and evaluating specialized, explainable LLMs for software security.
+Index Terms:
+Vulnerability Detection, Vulnerability Explanation, Large Language Model, Instruction Tuning.
 I Introduction
-A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2].
-These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation.
-The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
-Current vulnerability detection techniques mainly include pattern based and deep learning based approaches.
-Pattern based approaches [3, 4] generally rely on manually defined rules to detect vulnerabilities.
-Deep learning based approaches [1, 2, 5, 6, 7] train the models using existing vulnerability data and various code representation techniques.
-Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities.
-This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage.
-It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations.
-Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
-In the context of enhancing detection methods, the advent of Large Language Models (LLMs) [8, 9, 10] offers a promising avenue.
-With their advanced generative capabilities, LLMs have demonstrated significant potential across a wide range of applications, such as natural language processing and machine translation.
-These models can generate extensive textual content and provide contextually relevant information, making them well-suited for tasks that require detailed explanations.
-However, a considerable gap exists between the current capabilities of LLMs, particularly open-source models, and the specific requirements of vulnerability detection and explanation.
-Without domain-specific knowledge of vulnerabilities, LLMs struggle to effectively detect vulnerabilities and offer accurate explanations, which demand a deep understanding of code structures, security contexts, and the intricate interplay between various software components.
-To bridge this gap, it is essential to enhance LLMs with specialized capabilities for vulnerability detection and explanation.
-Fine-tuning open-source LLMs [11, 12, 13] on task-specific data presents an effective path toward achieving model specialization.
-Figure 1: Illustration of Vulnerability Detection Explanation.
-The main challenges in this specialization fine-tuning lie in two areas: (1) acquiring high-quality vulnerability explanation data [1, 14, 10, 15, 16], and (2) injecting vulnerability-related domain knowledge into LLMs [17].
-The improvement of model specialization depends heavily on high-quality, domain-specific data.
-However, in the field of vulnerability detection, there is a significant lack of annotated explanation data.
+Software vulnerabilities are flaws or weaknesses in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3], potentially leading to severe consequences such as data breaches, financial losses, and reputational damage.
+The growing complexity and interconnectedness of modern software systems pose significant challenges to the timely and effective identification and mitigation of these vulnerabilities.
+Vulnerability detection techniques are predominantly divided into pattern-based and deep learning-based approaches.
+Pattern-based methods [4, 5] rely on manually crafted rules, while deep learning-based approaches [1, 2, 6, 7, 8, 9, 10] leverage existing vulnerability datasets and various code representation techniques to train predictive models.
+Despite their advancements, a critical limitation of existing methods is their failure to provide detailed, human-understandable explanations for the vulnerabilities they detect.
+This explanatory gap severely impedes a developer’s ability to comprehend, validate, and effectively mitigate the identified risks in practice.
+Therefore, there is a pressing need for techniques that not only detect vulnerabilities but also provide insightful explanations, as illustrated in Figure 1.
+The figure contrasts a standard detection result with an enriched one, highlighting how comprehensive explanations aid in understanding and remediation.
+Figure 1: Illustration of Vulnerability Detection with and without Explanation. An explanation provides comprehensive information that aids in understanding and fixing the vulnerability.
+The advent of Large Language Models (LLMs) [11, 12, 13, 14, 15] offers a promising approach to bridging this explanatory gap.
+A prominent strategy to adapt these general-purpose models for specialized tasks is the “teacher-student” paradigm [16], where a “student” model is fine-tuned on data generated by a powerful “teacher” to distill its capabilities.
+However, applying this paradigm to the security-critical domain of vulnerability analysis is uniquely challenging.
+The goal is not merely to generate descriptive text, but to instill a model with the capacity for specialized, autonomous reasoning to both detect and explain vulnerabilities in real-world code.
+This requires a deep understanding of code semantics and security principles—a level of expertise far beyond general Q&A—presenting a formidable research challenge in transforming LLMs into specialized security assistants, which this work aims to address.
+The primary research challenges in this endeavor are twofold: (1) Acquiring High-Quality Vulnerability Explanation Data at Scale: In the field of vulnerability detection, there is a significant lack of annotated explanation data.
 Additionally, annotating such data typically requires costly manual labeling by domain experts, making the collection of large-scale annotated datasets a major challenge for model fine-tuning.
-Additionally, due to the inherent complexity of vulnerability detection and explanation tasks, merely providing task descriptions in prompts is insufficient for generating accurate answers.
-The model requires additional, critical vulnerability domain knowledge to identify and explain vulnerabilities accurately.
-To tackle these challenges, this study introduces LLMVulExp, an innovative automated framework designed to specialize open-source LLMs for software vulnerability tasks.
-Our framework directly addresses the two key challenges identified: the acquisition of high-quality annotated explanatory data and the effective injection of domain-specific vulnerability knowledge into LLMs.
-To address the first challenge, we leverage prompt-based techniques to automatically annotate explanatory information for vulnerability data.
-This annotated data includes both the vulnerability location and a detailed explanation, providing a rich dataset for fine-tuning.
-This approach reduces reliance on costly manual labeling by domain experts, enabling the collection of large-scale, annotated datasets necessary for model specialization.
-In addition, we propose new evaluation metrics for vulnerability explanations and an automated ’LLM-as-a-judge’ evaluation method [18, 19], ensuring a robust and scalable approach to verify the correctness and coherence of the generated explanations.
-For the second challenge, we apply instruction tuning [20] with Low-Rank Adaptation (LoRA) fine-tuning methods [12] to refine the open-source LLMs for injecting sufficient vulnerability patterns and knowledge.
-By focusing on task-specific adaptation, we guide the model to develop a deeper understanding of vulnerability semantics.
-To further enhance the model’s structural understanding of vulnerabilities, we adopt a Chain-of-Thought (CoT) strategy [21], guiding the model to concentrate on key code snippets most prone to vulnerabilities.
-Overall, the contributions of this paper are threefold:
-  • We introduce a comprehensive and effective workflow for training, inferring, and evaluating LLMs specifically on vulnerability detection and explanation tasks.
-  • We pioneer the exploration of the effectiveness of LLM-based vulnerability explanation generation and propose a tailored evaluation method and metrics for these explanations.
-  • We conduct extensive experiments to evaluate and analyze the feasibility of LLM-based vulnerability explanations, providing valuable insights for the practical application of specialized models in vulnerability detection.
+(2) Instilling Autonomous End-to-End Vulnerability Explanation Capability: The goal is to develop a “student” LLM that can autonomously analyze unfamiliar code, detect potential vulnerabilities, and generate accurate, actionable explanations—a task far more demanding than merely explaining pre-identified issues [17].
+Effectively fine-tuning a model for this integrated detection-explanation workflow, and determining if learning to explain concurrently enhances detection, remain key open questions.
+To address these challenges, we introduce LLMVulExp, a framework centered on a meticulously designed “teacher-student” paradigm that distinguishes between two LLM roles.
+First, we task a powerful “teacher” LLM with an “open-book” style problem: given the ground-truth vulnerability type and location, it generates a high-quality explanation articulating why the known flaw exists.
+This rich explanatory data is then used to train a specialized “student” LLM for the “closed-book,” real-world task of autonomously detecting and explaining vulnerabilities in unfamiliar code.
+Critically, this means the capability distilled from the teacher is not a magical detection ability, but rather the ability to structure and articulate complex security logic, from which the student develops its own autonomous reasoning and detection skills.
+To address the first challenge, our approach leverages powerful “teacher” LLMs.
+By providing them with ground-truth vulnerability type information, we guide them to generate high-quality explanatory data at scale on three vulnerability detection benchmark: SeVC [18], Diversevul [19] and PrimeVul [20].
+To ensure rigorous assessment, we also introduce a novel evaluation protocol for the generated explanations, comprising specific metrics—Accuracy, Clarity, and Actionability—and an automated “LLM-as-a-judge” methodology [21, 22], addressing a critical gap in evaluation standards.
+To address the second challenge, we employ a sophisticated fine-tuning process.
+Using the synthesized data, we specialize the “student” LLM for its autonomous, dual-capability role: first detecting vulnerabilities in unseen code and then explaining them.
+Our methodology explicitly investigates whether this explanation-focused training also enhances detection accuracy.
+We achieve this specialization efficiently using instruction tuning [23] with Low-Rank Adaptation (LoRA) [24] and promote deeper reasoning during inference via a Chain-of-Thought (CoT) prompting strategy [25].
+Overall, the primary contributions of this paper are threefold [^1] The replication package of our paper is available [26].:
+  • We introduce LLMVulExp, a novel “teacher-student” framework whose core technical innovation is a structured pipeline that leverages guided LLMs to resolve the critical data bottleneck in training explainable vulnerability detection models.
+  • We propose and validate a new evaluation protocol for vulnerability explanations, featuring dedicated metrics (Accuracy, Clarity, Actionability) and an “LLM-as-a-judge” method, thereby advancing the assessment standards for LLM-generated security analyses.
+  • Through extensive experiments, we demonstrate that our fine-tuning process equips “student” LLMs (e.g., CodeLlama) to autonomously detect vulnerabilities and generate accurate, clear, and actionable explanations.
+Our results provide valuable insights into the practical application of specialized models for end-to-end vulnerability detection and explanation.
 Paper Organization.
@@
 II-A Vulnerability Detection
-Prior studies proposed a series of deep learning approaches [1, 2, 5, 6, 7] to detect vulnerabilities.
+Prior studies proposed a series of deep learning approaches [1, 2, 6, 7, 8, 27, 28, 29] to detect vulnerabilities.
 These methods have employed labeled vulnerability data to train neural networks, enabling the models to capture semantic features associated with vulnerabilities.
@@
 II-B Vulnerability Explanation
-Although substantial progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these methods for vulnerability explanation remains a significant challenge.
-Despite the critical importance of reducing reliance on security experts and aiding developers in timely and thorough vulnerability mitigation, this area remains underexplored.
-Only a few studies focus on the explanatory capabilities of deep learning based models for vulnerabilities.
-VulDeeLocator [29] enhances a Bi-LSTM detector with an inner multiplication layer to forecast vulnerable statements.
-IVDetect [2] and LineVul [30] utilize subgraphs or attention weights from trained detectors to identify vulnerabilities.
-VELVET [31] integrates graph-based and sequence-based neural networks to prioritize vulnerable statements.
-LineVD [32] treats statement-level vulnerability detection as a node classification task, combining graph neural networks with a transformer-based model on the Program Dependency Graph(PDG).
-VulTeller [33] focuses on control and taint flows to detect dependencies for localizing vulnerabilities.
-VulExplainer [17] uses graph neural networks to locate fine-grained vulnerability information.
-While neural network-based models offer promising semantic understanding for vulnerability explanation, accurately detecting and explaining vulnerabilities using LLMs remains a substantial challenge for general-purpose code models.
-Our work explores the feasibility of fine-tuning specialized LLMs for both vulnerability detection and explanation, addressing a critical gap in the research literature.
+Vulnerability explanation in software security involves providing human-understandable insights into why a piece of code is vulnerable—encompassing its cause, potential exploitation, impact, and remediation—thereby extending beyond mere statement-level localization.
+Despite its critical importance for aiding developers and reducing reliance on security experts, achieving comprehensive explanations with deep learning models remains a significant, underexplored challenge.
+Prior works have made strides in aspects related to Vulnerability explanation.
+Many, such as VulDeeLocator [38], IVDetect [2], LineVul [39], VELVET [40], and LineVD [41], focus on statement-level vulnerability localization, pinpointing vulnerable code lines using techniques like Bi-LSTMs or graph-based analysis.
+While this localization is a crucial first step, the outputs often require considerable human interpretation to form a complete explanation.
+Other studies like VulTeller [10] and VulExplainer [17] offer deeper structural insights that contribute to detection.
+However, these approaches generally do not generate the complete, natural-language narratives necessary to fully explain the vulnerability’s nature and implications to developers.
+Accurately generating comprehensive and human-readable explanations remains a challenge for existing neural models.
+Our work explores fine-tuning specialized Large Language Models for both vulnerability detection and, crucially, for producing richer, more holistic explanations.
+We posit that LLMs, with their strong semantic understanding and text generation capabilities, can synthesize information to elucidate the ’why’ and ’how’ of vulnerabilities in coherent language.
+This aims to move beyond localization or basic structural cues, addressing a critical gap by providing actionable, understandable feedback directly to developers.
 III Methodology
@@
 In the context of vulnerability detection and explanation, it is crucial to ensure the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples for each type.
-In this paper, we conduct the study on two C/C++ function-level vulnerability datasets: (1) SeVC [34], which contains four core vulnerability types and over 50,000 vulnerable code snippets, and (2) DiverseVul [14], which covers 295 real open-source projects and 150 CWE types.
+In this paper, we conduct the study on three C/C++ function-level vulnerability datasets: (1) SeVC [18], which contains four core vulnerability types and over 50,000 vulnerable code snippets, (2) DiverseVul [19], which covers 295 real open-source projects and 150 CWE types and (3) PrimeVul [20], a new vulnerability dataset with high-quality, accurately labeled data and stringent de-duplication to offer realistic training and testing data for vulnerability detection.
 We select C/C++ as the programming language due to its diverse range of vulnerability types and examples, as well as the substantial dependence of fine-tuning on the availability of large-scale data.
 TABLE I: Statistics of the studied datasets.
-Dataset Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
-SeVC 56,395 40,491 4 Single/Multi-Type
-DiverseVul 18,945 9,161 10 Multi-Type(CWE)
+Dataset Non-Vul # Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
+SeVC 364,232 56,395 40,491 4 Single/Multi-Type
+DiverseVul 330,492 18,945 9,161 10 Multi-Label
+PrimeVul 228,800 6,968 4,041 10 Multi-Class
 Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types, comprising 56,395 vulnerable samples and 364,232 non-vulnerable ones.
@@
 It serves as a benchmark for evaluating the capability of LLMs in handling real-world scenarios with a wider variety of vulnerabilities.
-For our fine-tuning and evaluation target, we selected the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
-To ensure data quality, we deduplicate the vulnerability samples using the SHA-256 [35] hash method and downsample the non-vulnerable samples at a 1:1 ratio to match the vulnerable samples, achieving a balanced dataset, which aims to reduce training overhead and mitigates potential model bias.
-The details of the datasets are shown in Table I.
-We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerable and non-vulnerable samples to conduct our experiments.
+For our fine-tuning and evaluation target, we select the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
+PrimeVul is a state-of-the-art benchmark for C/C++ vulnerability detection, designed to overcome the critical data quality issues of prior datasets.
+It ensures high label accuracy (up to 92% ) through novel automated techniques, such as correlating code changes directly with official NVD descriptions and analyzing single-function security commits.
+This high-fidelity labeling, combined with stringent de-duplication and realistic chronological data splitting, presents a uniquely challenging testbed.
+Its difficulty is underscored by the original study, which shows that the performance of powerful models plummets on this benchmark—F1 scores drop from over 68% on older datasets to just 3.09%, and even GPT-4 fails to outperform random guessing on its challenging paired evaluation.
+Therefore, PrimeVul serves as a stringent benchmark to assess a model’s ability to generalize and reason about complex, real-world vulnerabilities, making it an ideal choice for validating our approach.
+We first deduplicate all samples using the SHA-256 hash method and then partition the data into training (80%), validation (10%), and test (10%) sets.
+To save training cost and mitigate model bias arising from severe class imbalance, we create balanced training and validation sets by downsampling non-vulnerable samples to a 1:1 ratio.
+The handling of the test set is tailored to the dataset’s characteristics.
+For the two datasets composed of real-world vulnerabilities, we preserve the original, imbalanced class distribution to ensure the evaluation reflects a realistic deployment scenario.
+In contrast, for the SeVC dataset, which contains a larger number of synthetic vulnerability samples, we use a balanced test set to facilitate a more effective multi-class evaluation.
+This overall approach allows for unbiased model training while rigorously testing its practical effectiveness.
+The final dataset statistics are detailed in Table I.
 ❷ Automated Vulnerability Explanation Annotation: Traditional open-source vulnerability datasets mainly contain source code, vulnerability labels, CWE types, and commit messages.
@@
 The model is guided by instruction-based prompt templates, paired with well-annotated examples, which stimulate its contextual learning capabilities, ensuring that the generated explanations are both accurate and informative.
-We implement this annotation process using GPT-3.5 [37], accessed via the API provided by OpenAI [38] due to its balance of annotation cost and efficiency.
-As part of our experimental study, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively.
+We implement this annotation process using LLM, selecting appropriate models based on a balance of cost, efficiency, and capability.
+For our initial annotation of the SeVC and DiverseVul datasets, we employed GPT-3.5 [43], accessed via the API provided by OpenAI [44].
+For the subsequently added experiments and the PrimeVul dataset, we utilized Deepseek-V3 [45], a state-of-the-art model noted for its strong code processing capabilities.
+This dual-model approach allowed us to complete a large-scale annotation effort, resulting in 40,491, 9,161 and 4,041 vulnerability explanation data points across the respective datasets.
 This effort fills a significant gap in the availability of vulnerability explanation annotations, enabling more effective training and evaluation of vulnerability detection and explanation models.
@@
 This approach enables the model to focus on the most relevant parts of the code, thereby improving accuracy and interpretability.
-Using the prompt template illustrated in Figure LABEL:fig:prompt_key, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type.
+Using the prompt template illustrated in Figure 4, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type.
 These extracted statements form a structured CoT, guiding the fine-tuned model through a step-by-step process for targeted vulnerability detection, location, and explanation.
@@
   • RQ3: How do explanations affect the results of vulnerability detection?
-  • RQ4: How does the key code extraction impact detection performance?
+  • RQ4: How does the key code extraction impact detection and explanation performance?
 IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
-Accurate detection is the foundation for correct vulnerability explanation.
-In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
+Accurate detection is the foundation for any subsequent vulnerability analysis or explanation.
+In this RQ, we first conduct a comprehensive evaluation of our fine-tuned model’s detection performance against strong baselines across two distinct, real-world scenarios.
 IV-A1 Experimental Setup.
-Dataset.
-As discussed in Section III, we select the SeVC [34] dataset and the DiverseVul dataset [14].
-Based on the characteristics of the two datasets, we constructed three different vulnerability detection tasks to evaluate the detection performance of LLMVulExp: binary classification vulnerability detection (SeVC), coarse-grained multi-class vulnerability detection (SeVC), and CWE type-based multi-label vulnerability detection (DiverseVul).
-Backbone LLMs.
-In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [39] as our primary model for the experiments discussed in this chapter.
-CodeLlama is initialized with the weights of Llama2 [8] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities.
-Experimental Setting.
-To comprehensively evaluate the fine-tuned vulnerability LLMs, we design three sub-research questions:
-  • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
-  • (RQ1.2) How effective is the model in detecting multiple types of vulnerabilities simultaneously?
-  • (RQ1.3) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
-To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics.
-For the binary classification scenario in RQ1.1, we directly use these three metrics.
-For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1.
-For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1.
-To test detection accuracy, we selected CodeT5 [40] and CodeBERT [41] as baselines and fine-tuned them for classification tasks by adding a linear classification layer.
-These models are chosen for two primary reasons: (1) As pre-trained code models, CodeT5 and CodeBERT have demonstrated consistent and superior performance on prior vulnerability detection benchmarks [16, 14], making them reliable baselines.
-(2) Their inclusion allows us to assess whether fine-tuned LLMs offer an advantage over existing pre-trained code models in terms of detection accuracy.
-It is important to note that the focus of this study is not on developing the most optimal detection model but rather on evaluating the explanation capabilities of fine-tuned LLMs.
-Therefore, using these two baselines provides a meaningful comparison without detracting from the primary objective of assessing the interpretability of LLM-generated explanations.
-Implementation Details.
-We implement our approach using the Transformers[13] and PEFT[42] libraries on the PyTorch platform.
-All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs, with a token length limit set to 2048.
-The models are trained using the AdamW optimizer for 3 epochs with a batch size of 2 and data precision set to float16 to optimize memory usage and computational efficiency.
-We set the learning rate to 0.0003 and weight decay to 0.01.
-For the LoRA configuration, we set target_modules to {q_proj, k_proj, v_proj, o_proj} for SeVC and {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj} for DiverseVul.
-The LoRA rank is set to 16, the LoRA scaling factor to 16, and the dropout rate to 0.05.
-For fine-tuning CodeBERT and CodeT5, we use the AdamW optimizer with the learning rate of 1e-5 and the weight decay of 1e-2, and train the models for 3 epochs.
-All other parameters are set to their default values.
+Datasets and Tasks.
+As detailed in Section III, we select three benchmarks for evaluation: SeVC [18], DiverseVul [19] and PrimeVul [20].
+Based on their distinct characteristics, we formulate corresponding detection tasks:
+  • On SeVC, we peform binary classification vulnerability toward single-type vulnerability.
+  • On SeVC, we perform coarse-grained multi-class vulnerability detection.
+  • On DiverseVul, we perform multi-label CWE classification (a function can have multiple vulnerability types) and binary classification.
+  • On PrimeVul, we perform multi-class CWE classification (a function has exactly one vulnerability type) and binary classification.
+Model and Baselines.
+Our primary model, LLMVulExp, is based on CodeLlama-13B-Instruct [46], fine-tuned using the LoRA technique.
+To rigorously situate its performance, we compare it against a comprehensive set of four baselines: two general-purpose pre-trained models, CodeT5 [47] and CodeBERT [48], and two state-of-the-art specialized vulnerability detectors, LineVul [39] and CausalVul [49].
+While our ultimate goal includes evaluating explanation capabilities, validating detection efficacy against these strong baselines is a critical first step.
+Metrics and Implementation Details.
+We use Precision, Recall, and F1-Score as our primary evaluation metrics.
+For the multi-label task on DiverseVul, we report Micro-F1 and Macro-F1.
+For the multi-class task on SeVC and PrimeVul, we report Weighted-F1 and Macro-F1 to analyze performance across its imbalanced class distribution.
+We implement our approach using the Transformers [50] and PEFT [51] libraries.
+All experiments are conducted on two NVIDIA A100-80GB GPUs.
+Key hyperparameters include a learning rate of 3e-4, 3 training epochs, and a batch size of 2.
+For the LoRA configuration, we set {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj}.
+The LoRA rank is set to 16, the LoRA scaling factor to 16/32, and the dropout rate to 0.05.
 IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model.
@@
 This highlights the critical role of balanced and representative training datasets in improving the model’s performance across all vulnerability types.
-IV-A4 RQ1.3: Identification of a greater variety of vulnerability types.
-In real-world software development environments, security risks arise from a diverse range of project types and complex code structures, often involving a broader spectrum of vulnerability categories.
-To better reflect these real-world conditions, we conduct a multi-label classification task using the top 10 CWE categories in the real-world vulnerability dataset DiverseVul, where a subset of the code samples is annotated with multiple CWE labels.
-Specifically, the task description is modified to emphasize CWE-type vulnerability detection and explanation, with the addition of a ‘[CWE]’ tag to generate a list of CWE categories relevant to the target code.
-Furthermore, CWE descriptions are incorporated as part of the output explanation, aiding the model in comprehending the context and semantics of each CWE type and leveraging these descriptions for vulnerability analysis and identification.
-TABLE V: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.3).
-Metric CWE ID Overall (Micro-F1)
-787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
-Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 70.8% 65.3%
-Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 63.1% 61.9%
-F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 67.0% 63.5%
-Experimental Results.
-The experimental results are presented in Table V.
-The findings demonstrate that after fine-tuning, the specialized vulnerability models achieved high detection accuracy across the 10 CWE categories, highlighting their capability to capture the distinctive patterns associated with each CWE type.
-Unlike the multi-class task in SeVC, the multi-label classification in DiverseVul does not require the model to distinguish between mutually exclusive categories.
-Instead, each CWE type has a more granular and explicit definition, mitigating the category association challenges observed in SeVC and thereby yielding improved performance.
-These results validate the practicality of employing LLMs for vulnerability detection and explanation in real-world software development scenarios.
-The following example of a generated explanation further illustrates the model’s capacity to analyze and identify CWE-based vulnerabilities effectively.
-1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.
+IV-A4 RQ1.3: Performance on the DiverseVul Dataset
+Task Setup.
+To evaluate our model’s effectiveness in a standard, realistic scenario, we utilize the DiverseVul dataset.
+This benchmark is particularly suitable as it reflects real-world conditions where code can contain multiple, distinct types of vulnerabilities.
+We assess the models on two fronts: (1) a multi-label classification task on the top 10 most frequent CWE types, and (2) a standard binary classification task.
+Experimental Results. The comprehensive results on DiverseVul are presented in Table V.
+On the binary detection task, a notable observation is that all models achieve a perfect Precision of 100.0%.
+This suggests the non-vulnerable samples in the test set are sufficiently distinct, leading to a near-zero false positive rate.
+Therefore, the key differentiating metric is Recall.
+Here, LLMVulExp demonstrates a decisive advantage, achieving a Recall of 97.7%, which significantly surpasses all baselines by over 34 percentage points.
+This superiority results in a state-of-the-art F1 Score of 98.8%.
+In the more challenging multi-label CWE detection task, our model continues to show robust performance.
+It achieves the highest Macro-F1 69.9% and Micro-F1 70.5% scores, indicating the most balanced and effective performance overall.
+This success is primarily driven by its outstanding Macro-Recall and Micro-Recall, once again highlighting its strength in comprehensively identifying all relevant vulnerabilities.
+While some baselines like CodeT5 attain slightly higher Precision, their significantly lower Recall cripples their overall effectiveness.
+These strong results validate the capability of our fine-tuned LLM approach for nuanced vulnerability detection.
+The example in Figure IV-A4 further illustrates our model’s capacity to generate accurate and context-aware explanations.
+TABLE V: Performance Comparison on DiverseVul (RQ1.3).
+All metric values are percentages (%).
+Best results in each row are in bold.
+* means indicates statistically significant improvement over the baseline method through paired t-test ($p<0.05$).
+Task Metric Ours CodeBERT CodeT5 LineVul CausalVul
+Multi-Label CWE Detection Precision (Micro) 71.7 65.3 72.8 65.0 67.3
+Recall (Micro) 69.3 38.2 56.8 38.1 37.4
+F1 Score (Micro) 70.5 48.0 63.9 48.0 48.3
+Precision (Macro) 72.5 69.0 73.5 68.9 69.0
+Recall (Macro) 67.8 32.0 52.7 32.3 32.2
+F1 Score (Macro) $\textbf{69.9}^{*}$ 39.0 60.6 40.2 39.1
+Binary Vulnerability Detection Precision 100.0 100.0 100.0 100.0 100.0
+Recall 97.7 61.1 62.9 63.2 60.7
+F1 Score $\textbf{98.8}^{*}$ 75.9 77.2 77.4 75.5
+1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 }
 [cwe] This function is related to [’CWE-476’].
@@
 TABLE VI: Example of Generated Explanation on DiverseVul.
-Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
+IV-A5 RQ1.4: Performance on the PrimeVul Dataset
+Task Setup.
+To further probe the limits of current models, we conduct a more stringent evaluation using the PrimeVul dataset.
+We This task is formulated as multi-class classification, where each vulnerable function belongs to exactly one CWE type.
+PrimeVul is recognized as a more challenging benchmark due to its wider variety of vulnerability types and more realistic, imbalanced data distribution, which rigorously tests a model’s ability to distinguish between fine-grained vulnerability patterns.
+Experimental Results. The performance on PrimeVul, detailed in Table VII, presents a complex and challenging picture that highlights the critical issue of class imbalance in vulnerability detection.
+In the multi-class CWE detection task, the high Weighted-F1 scores (e.g., 94.9% for LLMVulExp) initially suggest strong performance.
+However, these scores are misleading as they are dominated by the models’ accuracy on the major non-vulnerable code.
+The performance of all models plummets on the Macro-averaged metrics, with the top Macro-F1 score being a mere 21.7% from CausalVul, and our model achieving a comparable 20.7%.
+This dramatic gap between Weighted and Macro scores proves that all models, while able to learn common vulnerability patterns, largely fail to generalize to the long tail of less frequent CWEs—a critical failure for real-world applications.
+This difficulty is further confirmed by the binary detection task.
+On this more challenging dataset, all models experienced a drastic performance drop compared to their results on DiverseVul.
+Our model, LLMVulExp, while achieving the highest F1-Score at 39.8%, still operates at a level far below what is required for reliable, practical use.
+The findings from PrimeVul strongly suggest that current fine-tuning methodologies for LLMs, despite their promise, are not yet robust enough for complex, imbalanced, real-world vulnerability analysis.
+This underscores a critical research gap and motivates the need for more advanced techniques to enhance model generalization and robustness against low-frequency data.
+TABLE VII: Performance Comparison on PrimeVul (RQ1.4).
+All metric values are percentages (%).
+Best results in each row are in bold.
+* means indicates statistically significant improvement over the baseline method through paired t-test ($p<0.05$).
+Task Metric Ours CodeBERT CodeT5 LineVul CausalVul
+Multi-Class CWE Detection Precision (Weighted) 96.2 96.0 95.9 95.2 96.1
+Recall (Weighted) 93.1 90.2 89.2 93.4 90.4
+F1 Score (Weighted) 94.9 93.1 91.9 94.1 93.3
+Precision (Macro) 19.2 13.7 18.7 14.4 25.0
+Recall (Macro) 25.7 20.1 20.4 19.3 19.2
+F1 Score (Macro) 20.7 14.1 15.3 14.0 21.7
+Binary Vulnerability Detection Precision 29.5 19.6 18.2 25.8 20.2
+Recall 61.1 60.1 61.1 47.8 59.4
+F1 Score $\textbf{39.8}^{*}$ 29.6 28.0 33.5 30.1
+Summary of RQ1: Our evaluation offers a dual perspective on LLMs for vulnerability detection.
+It confirms their feasibility and effectiveness on established benchmarks like SeVC and DiverseVul.
+However, it also reveals their current limitations, as all models struggled with the class imbalance and complexity of the more challenging PrimeVul dataset, highlighting a critical area for future research.
 IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
@@
 The explanatory content generated by the model should assist software developers in identifying and mitigating potential vulnerabilities by ensuring the accuracy of the analysis, maintaining readability and conciseness, and providing actionable suggestions for remediation.
-The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a ”correct” or ”complete” explanation.
+The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a “correct” or “complete” explanation.
 This inherent subjectivity poses significant challenges for effective evaluation.
@@
 IV-B1 Experimental Setup.
-In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the SeVC and DiverseVul test sets, alongside ground truth annotations generated using GPT-3.5.
+In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the DiverseVul and PrimeVul test sets, alongside ground truth annotations generated using Deepseek-V3.
 Given the large scale of the dataset and the high cost of evaluation, we employ a random sampling approach based on a 95% confidence level and a 5% confidence interval [54].
-This yields 384 and 224 samples from SeVC and DiverseVul, respectively, ensuring statistical validity and representativeness in our evaluation.
+This yields 224 and 197 samples from DiverseVul and PrimeVul, respectively, ensuring statistical validity and representativeness in our evaluation.
 Our evaluation includes two parts: (1) Manual review: Two authors, with extensive research experience in software vulnerabilities and development, independently review the sampled data, scoring them based on predefined criteria.
 Disagreements are resolved through discussion until a consensus is reached.
-The Cohen’s Kappa value [46] for inter-reviewer agreement is 0.76, indicating substantial consistency and reliability.
-(2) Automated review using LLM: GPT-3.5 is used for automated evaluation.
+The Cohen’s Kappa values [55] for inter-reviewer agreement are 0.78 and 0.81 for two datasets, indicating substantial consistency and reliability.
+(2) Automated review using LLM: Deepseek-V3 is used for automated evaluation.
 A detailed prompt is provided, describing the vulnerability explanation evaluation task and the specific criteria.
@@
 IV-B2 Experimental Results.
-We present the results of our dual evaluation process in Table VII.
-The suffix Gen.
-refers to explanations generated by LLMVulExp, while Ann.
-refers to annotations from GPT-3.5.
+We present the results of our dual-rater (Manual and LLM-Automated) evaluation in Table VIII.
+The column suffixes denote the data source: -Gen refers to explanations generated by our model (LLMVulExp), while -Ann refers to the higher-quality explanations produced by our annotation method.
 All-Pos.
-represents cases meeting all three evaluation metrics.
-Key findings include:
-(1) Accuracy of Explanations: LLMVulExp achieves high accuracy (e.g., over 90.0% for SeVC-Gen.
-in both manual and automated reviews), pinpointing vulnerability risks and code locations.
-This underscores the importance of accurate vulnerability type identification.
-(2) Clarity of Explanations: The clarity scores are 81.4% for SeVC and 94.1% for DiverseVul, indicating the model effectively reduces cognitive barriers for developers analyzing vulnerabilities.
-(3) Actionability of Explanations: Explanations from LLMVulExp include actionable suggestions for code modifications, achieving 93.4% for SeVC and 80.5% for DiverseVul.
-This highlights the model’s potential to support practical remediation efforts.
-(4) Effectiveness of Annotation Method: Results from SeVC-Ann.
-and DIV-Ann.
-validate the effectiveness of our annotation method in generating high-quality explanatory data, reducing subsequent annotation costs.
-(5) Potential of LLM in Automated Assessment: LLM-based automated evaluations closely align with manual reviews.
-For All-Pos., manual vs.
-automated results are 76.0% vs.
-74.0% for SeVC and 80.1% vs.
-71.9% for DiverseVul.
-This demonstrates LLMs’ potential to scale evaluation tasks efficiently.
+is a strict metric representing cases that satisfy all three criteria (Accuracy, Clarity, and Actionability).
+Our analysis yields three key findings:
+(1) Generated Explanations Show Value, but Annotated Explanations Are Superior.
+Our model-generated explanations (-Gen) demonstrate practical value, achieving, for instance, a manual All-Positive score of 51.3% on DiverseVul.
+However, the results consistently show that the explanations produced via our annotation method (-Ann) are of significantly higher quality across all metrics and datasets.
+For example, on PrimeVul, the manually-assessed All-Positive score for PRI-Ann (90.3%) is more than double that of PRI-Gen (41.6%).
+This validates the effectiveness of our annotation pipeline in creating a high-quality dataset for fine-tuning and evaluation.
+The lower scores on PrimeVul compared to DiverseVul also echo our findings from RQ1, confirming it as a more challenging benchmark for both detection and explanation.
+(2) Agreement Varies Significantly Between Objective and Subjective Metrics.
+The Cohen’s Kappa coefficients, which measure the agreement between manual and LLM raters, reveal a critical insight.
+For the relatively objective Accuracy metric, which assesses factual correctness (e.g., correct location, valid CWE), the agreement is consistently in the moderate to substantial range (Kappa from 0.52 to 0.65).
+This indicates that LLMs can reliably replicate human judgment when the evaluation criteria are concrete and verifiable.
+In stark contrast, the agreement for the more subjective metrics of Clarity and Actionability is significantly lower and more varied.
+On DiverseVul, the Kappa values are near zero (0.01 to 0.12), signaling almost no agreement beyond chance.
+While the agreement is better on PrimeVul, it remains modest.
+This discrepancy strongly suggests that while LLMs are proficient at automating factual checks, they struggle to consistently align with human assessment of subjective qualities like writing style, comprehensibility, and the perceived usefulness of repair suggestions.
+(3) LLMs Are Viable for Automating Concrete, Fact-Based Evaluation.
+Based on the Kappa analysis, we conclude that LLMs are most suitable for automating the evaluation of specific, clearly-defined aspects of an explanation’s quality.
+Our comparative analysis of inter-rater agreement reveals a clear and practical boundary for using LLMs in evaluation.
+Compared to human-human agreement, the human-LLM agreement is relatively lower.
+Crucially, this significant gap is not uniform.
+The discrepancy is primarily driven by the LLM’s struggle with nuanced, subjective metrics like Clarity and Actionability.
+The strong agreement on the ‘Accuracy‘ metric validates the potential of LLM-based automation for reliably checking the factual correctness of generated explanations at scale.
+This allows human experts to focus their limited time on the more nuanced, subjective aspects that still require human oversight.
+Our dual-validation strategy, therefore, offers a practical framework: leveraging LLMs for scalable, objective assessment while using targeted human review for subjective quality control, creating an efficient and reliable evaluation pipeline.
 While the proposed approach demonstrates strong performance, it has inherent limitations.
@@
 Despite these challenges, the dual-validation strategy combining manual and automated reviews offers a practical and effective solution for evaluating vulnerability explanations, balancing accuracy with scalability and cost-effectiveness.
-TABLE VII: Proportion (%) of Vulnerability Explanation Review Results (RQ2).
-Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
+TABLE VIII: Proportion (%) of Vulnerability Explanation Review Results and Inter-Rater Agreement (RQ2).
+Metric DIV-Gen. DIV-Ann. PRI-Gen. PRI-Ann.
 Manual
-Accuracy 91.1 93.1 73.3 94.1
-Clarity 81.4 81.7 94.1 98.6
-Construct. 93.4 94.6 80.5 83.2
-All-Pos. 76.0 76.0 59.7 80.1
+Accuracy 57.4 84.3 43.8 93.0
+Clarity 96.7 99.0 71.4 97.8
+Action. 82.7 83.2 42.7 90.3
+All-Pos. 51.3 71.1 41.6 90.3
 LLM-Automation
-Accuracy 90.4 97.6 96.4 96.8
-Clarity 74.3 77.2 95.0 96.8
-Construct. 83.5 88.6 67.9 71.9
-All-Pos. 72.8 74.0 67.9 71.9
-1int i; 2for(i = 0; i < 1; i++) 3twoIntsStruct * dataBuffer = (twoIntsStruct *) ALLOCA(100 * sizeof(twoIntsStruct)); 4size_t i; 5for (i = 0; i < 100; i++) 6 dataBuffer[i].intOne = 1; 7 dataBuffer[i].intTwo = 1; 8data = dataBuffer; 9printStructLine(&data[0]); 10 11void printStructLine(const twoIntsStruct * structTwoIntsStruct) { 12 printf($d\n$", structTwoIntsStruct->intOne, structTwoIntsStruct->intTwo); 13} 14delete [] data;
-{mdframed}[style=mystyle] [type] pointer
-[location] The line ”delete [] data” has a pointer-related issue.
-[explanation]
-(Analysis:) The issue in this code is related to the ”data” pointer, which is being improperly freed using the ”delete” operator.
-The ”data” pointer is assigned the address of a memory block allocated on the stack using the ”ALLOCA” function, which allocates memory on the stack instead of the heap.
-The ”delete” operator is used to free memory allocated on the heap using the ”new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
-(Suggestion:) To fix this issue, the program should use the ”free” function to free memory allocated on the stack, or use the ”new” operator to allocate memory on the heap instead of the ”ALLOCA” function.
-Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
-TABLE VIII: Example of Generated Explanation on SeVC.
-IV-B3 Illustrative Example.
+Accuracy 62.9 89.3 46.5 94.6
+Clarity 90.9 97.5 73.5 99.5
+Action. 61.4 85.3 47.0 93.5
+All-Pos. 59.9 85.3 46.5 93.0
+Agreement (Cohen’s Kappa)
+Human vs. Human Agreement (per-metric)
+Accuracy 0.92 0.90 0.77 0.75
+Clarity 0.76 0.66 0.82 0.85
+Actionability 0.60 0.60 0.64 0.65
+Human vs. LLM Agreement (per-metric)
+Accuracy 0.65 0.52 0.60 0.58
+Clarity 0.04 0.01 0.57 0.39
+Actionability 0.02 0.12 0.54 0.57
+Overall Agreement Comparison
+Human vs. Human 0.78 0.81
+Human vs. LLM 0.38 0.68
+IV-B3 Qualitative Analysis of Vulnerability Explanations
+Beyond the numerical scores, a qualitative analysis of the generated explanations is crucial to understand why the models perform as they do.
+By examining both exemplary and unsatisfactory cases, we can uncover the models’ underlying capabilities and current limitations.
+An Exemplary Case.
 To highlight the importance of evaluating vulnerability explanations, we analyze an example from SeVC.
 The sample pertains to a pointer-type vulnerability caused by incorrect memory deallocation in C code.
-As shown in Figure IV-B2, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function.
+As shown in Figure IV-B3, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function.
 The explanation provides actionable suggestions, such as ensuring pointer safety by using the appropriate deallocation methods.
 It is clear, logically structured, and actionable, satisfying all three evaluation criteria.
-Despite these strengths, some generated explanations exhibit issues, such as vague locations, verbose descriptions, or lack of actionable suggestions.
-While rare, hallucinations may occur, typically involving memorized examples or misinterpreted contexts.
-Overall, the fine-tuned model demonstrates robust capabilities for generating high-quality vulnerability explanations.
-Summary of RQ2: We find that LLMVulExp is capable of generating vulnerability explanations with high levels of Accuracy, Clarity, and Actionability.
-Additionally, we explore the potential of leveraging LLMs for automated data annotation, significantly reducing the reliance on manual effort.
+[location] The line “delete [] data” has a pointer-related issue.
+[explanation]
+(Analysis:) The issue in this code is related to the “data” pointer, which is being improperly freed using the “delete” operator.
+The “data” pointer is assigned the address of a memory block allocated on the stack using the “ALLOCA” function, which allocates memory on the stack instead of the heap.
+The “delete” operator is used to free memory allocated on the heap using the “new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
+(Suggestion:) To fix this issue, the program should use the “free” function to free memory allocated on the stack, or use the “new” operator to allocate memory on the heap instead of the “ALLOCA” function.
+Additionally, the program should ensure that the “data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the “data” pointer behaves correctly and safely in all cases.
+TABLE IX: Example of Generated Explanation on SeVC.
+Insights from Unsatisfactory Explanations.
+An analysis of the less satisfactory explanations reveals recurring patterns that not only explain the numerical results in Table VIII but also offer crucial insights for future research.
+1.
+Factual Errors vs.
+Subjective Deficiencies: Explaining the Kappa Discrepancy.
+The shortcomings in generated explanations can be broadly categorized into two types, which directly explain the variance in inter-rater agreement (Kappa scores).
+Factual Inaccuracies: The first category includes objective, verifiable errors.
+These are the primary reason for low Accuracy scores and include: (a) Incorrect Localization, where the model identifies the wrong line number; (b) Root Cause Misidentification, where it describes a symptom (e.g., a crash) instead of the underlying flaw (e.g., a use-after-free); and (c) Vulnerability Misclassification.
+Because these errors are factual, both human and LLM raters can assess them with high consistency, leading to the moderate-to-substantial Kappa agreement (0.52-0.65) observed for the Accuracy metric.
+Subjective Deficiencies: The second category involves subjective qualities of the explanation.
+These include: (a) Generic or Missing Repair Guidance, where suggestions are too vague (e.g., “validate all inputs”); and (b) Poor Clarity, due to excessive verbosity or templated, repetitive phrasing.
+Assessing these qualities is inherently subjective—what one developer finds clear, another may find verbose.
+This subjectivity is the primary reason for the low-to-minimal Kappa agreement seen for Clarity and Actionability, especially on the DiverseVul dataset.
+2.
+Shallow Pattern Matching vs.
+Deep Semantic Reasoning: The Core Limitation.
+A deeper insight is that many of these failures stem from a single core limitation: the model often relies on shallow pattern matching rather than deep, causal reasoning about the code’s execution flow.
+This hypothesis explains several observed phenomena.
+Incorrect Localization often results from matching superficial tokens (e.g., a function name) without understanding the control flow that activates the vulnerability.
+The model’s tendency to produce templated or generic explanations suggests it is retrieving and adapting a learned pattern rather than reasoning from first principles.
+Most importantly, this explains why performance was notably worse on the more complex PrimeVul dataset (Table VIII).
+The vulnerabilities in PrimeVul likely require a more global understanding of inter-procedural dependencies, which is a known weakness of models that rely on local, pattern-based reasoning.
+The model excels at explaining common vulnerabilities it has seen frequently during training but struggles when faced with novelty or complexity.
+Implications.
+The qualitative analysis reveals the dual nature of current LLMs for vulnerability explanation.
+They are powerful pattern recognizers, capable of producing excellent explanations for common vulnerabilities in relatively simple code.
+However, their reliability diminishes when faced with complex or novel scenarios that require deep semantic understanding.
+This underscores that while LLMs are invaluable assistive tools, they are not yet infallible replacements for human expertise.
+The path forward for research is clear: we must develop methods that move beyond pattern matching to instill a more robust, causal reasoning capability in these models.
+Summary of RQ2: Our analysis reveals a critical insight for automated explanation evaluation: LLMs show strong agreement with human experts on factual Accuracy but diverge significantly on subjective metrics like Clarity and Actionability.
+This suggests that the most promising path forward is a hybrid evaluation strategy, where LLMs automate scalable, objective checks, while human oversight remains essential for nuanced, subjective quality assessment.
 IV-C RQ3: How do explanations affect the results of vulnerability detection?
-Building on the detection capabilities evaluated in RQ1, we consider the detection task as a foundational step in the broader explanation task.
-This raises an important question: How does fine-tuning for vulnerability explanation influence the detection performance of LLMs compared to models fine-tuned solely for detection?
-In this RQ, we experimentally examine the impact of incorporating explanatory information on vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory data.
+Building on the detection capabilities evaluated in RQ1, this section investigates a critical question: Does the act of generating explanations help or hinder the primary task of vulnerability detection?
+To answer this, we experimentally examine the impact of the explanation task by comparing our standard model, which is fine-tuned to both detect and explain, against a variant fine-tuned solely for detection.
 Experimental Setup.
-We conduct ablation studies under the multi-type vulnerability detection scenario in SeVC (RQ1.2) and the multi-label CWE detection scenario in DiverseVul (RQ1.3).
-To isolate the effect of explanatory information, we transform the vulnerability explanation fine-tuning task into a detection-only task by removing explanatory data from annotations and adjusting task descriptions and prompts.
-The resulting LLMs output only vulnerability types, creating a detection-focused model (referred to as Detector).
-This is compared against the performance of the original explanation-focused model (referred to as Explainer).
-Figure 5: Performance of Explainer and Detector.
+We conduct this ablation study on the SeVC, DiverseVul and PrimeVul datasets.
+To isolate the effect of the explanation task, we create a detection-focused counterpart for our model.
+This is achieved by removing all explanatory text from the training data and adjusting the instruction prompts so that the model’s only goal is to output the vulnerability label(s).
+We refer to our standard model as the Explainer and its detection-only counterpart as the Detector.
+All other hyperparameters and training procedures are kept identical.
+Figure 5: Performance comparison of the Explainer (trained to explain and detect) and the Detector (trained only to detect) on the DiverseVul and PrimeVul datasets.
 Experimental Results.
-The comparative results between Explainer and Detector are illustrated in Figure 5, where ”Mi” and ”Ma” denote Micro and Macro metrics, respectively.
-Incorporating explanatory data into the fine-tuning process did not significantly reduce detection performance.
-For instance, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while both achieve 74.8% on DiverseVul.
-Interestingly, certain metrics improved when explanatory data was included, such as the Macro-F1 score on DiverseVul, where the Explainer outperformed the Detector by nearly 2% (73.2% vs.
-71.4%).
+We first show the comparison of Detector and Explainer on easier SeVC for multi-type vulnerability detection in Table X.
+Our experimental results reveal that detection-only models outperform interpretable models on the SeVC synthetic dataset.
+This occurs because synthetic vulnerability patterns in SeVC are highly discernible, and interpretation task training directs the model’s focus toward explanation generation, slightly compromising detection performance.
+The marginal performance discrepancy indicates strong compatibility between these tasks, demonstrating that interpretation tasks can jointly participate in large model fine-tuning alongside detection objectives.”
+TABLE X: SeVC Dataset Experiment Results Comparison
+Model Mi-Pre. Mi-Rec. Mi-F1 Ma-Pre. Ma-Rec. Ma-F1
+SeVC-Explainer 79.9 80.5 80.1 70.0 72.4 70.8
+SeVC-Detector 81.7 81.4 81.2 70.6 75.4 72.0
+The comparative results of real-world vulnerability detection, shown in Figure 5, reveal a nuanced and dataset-dependent relationship between explanation and detection performance.
+The impact of learning to explain is not monolithic; instead, it is dictated by the complexity of the underlying task.
+On the simpler benchmark (DiverseVul), the detection-only model (‘Detector‘) shows a marginal performance advantage.
+For instance, the ‘Detector‘ achieves a slightly higher Macro-F1 score than the ‘Explainer‘ (71.0% vs.
+69.9%).
+This suggests that on less complex tasks, forcing the model to also generate explanations may introduce a minor multi-task learning overhead, slightly diluting its focus on the primary detection objective.
+On the more challenging benchmark (PrimeVul), this trend dramatically reverses.
+Here, the ‘Explainer‘ model significantly outperforms the ‘Detector‘ across nearly all metrics.
+The most notable advantage is seen in the Micro-F1 score (93.2% vs.
+88.7%), but the trend also holds for the more challenging Macro-F1 score (20.7% vs.
+18.2%).
+This reversal is a critical finding.
+It strongly suggests that for complex, real-world vulnerabilities with subtle patterns, the task of generating explanations is not a hindrance.
+Instead, it acts as a form of implicit regularization or deep supervision.
+By learning to articulate why a piece of code is vulnerable, the model is compelled to develop a more robust and granular understanding of the vulnerability’s root cause.
+This deeper semantic comprehension, forged through the explanation task, directly enhances its ability to detect those same complex vulnerabilities, providing a performance benefit that far outweighs any multi-task overhead.
 These findings demonstrate that the vulnerability explanation task can coexist with the detection task without compromising the model’s detection accuracy.
@@
 In some cases, well-annotated explanatory data can even improve the model’s overall understanding and effectiveness across both tasks.
-Summary of RQ3: Integrating vulnerability explanations into the fine-tuning process preserves detection capabilities and may even enhance performance in specific vulnerability scenarios.
-IV-D RQ4: How does the key code extraction impact detection performance?
-In this RQ, we investigate whether LLMs can identify key code segments prone to vulnerabilities and enhance detection by focusing on these critical sections.
-To explore this, we annotate key code segments in the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the annotated key code information.
+Summary of RQ3: The relationship between explanation and detection is nuanced and depends on task complexity.
+For simpler vulnerabilities, a detection-only focus is marginally better.
+For complex, realistic vulnerabilities, however, fine-tuning for explanation significantly boosts detection performance by fostering a deeper, more robust understanding of root causes.
+IV-D RQ4: How does the key code extraction impact detection and explanation performance?
+In this RQ, we investigate whether explicitly guiding an LLM to focus on critical code segments can enhance its performance.
+To explore this, we creat datasets for SeVC, DiverseVul and PrimeVul where likely vulnerable code snippets (“key code”) are pre-annotated.
+We then fine-tune and evaluate our model using this focused information.
 Experimental Setup.
-As outlined in Section III, we employ GPT-3.5 to annotate key code segments closely related to vulnerabilities.
-To avoid label leakage, vulnerability-type tags are concealed during the annotation process.
-For training samples where key code information was unavailable, the annotations are nullified.
-Similarly, test samples without key code annotations are excluded from the evaluation.
-During fine-tuning and inference, the prompts are modified to emphasize task descriptions and instructions, guiding the model to focus on the extracted key code for vulnerability detection.
-TABLE IX: F1 of Key Code Extract. on SeVC: (RQ4).
-Type Single-Type(RQ1.1) Multi-Type(RQ1.3)
+As outlined in Section III, we employ LLMs to annotate key code segments related to vulnerabilities in our datasets.
+To prevent data leakage, vulnerability type information is concealed during this annotation process.
+During the fine-tuning and inference stages for the “With Key Code” model, our prompts are modified to instruct the model to specifically leverage these extracted snippets for its analysis.
+We then compare this model’s detection performance against the baseline model trained “Without Key Code”.
+We also briefly assess the impact on explanation quality using the same methodology as in RQ2.
+TABLE XI: F1 of Key Code Extract. on SeVC: (RQ4).
+Type Single-Type(RQ1.1) Multi-Type(RQ1.2)
 W/O Key. With Key. W/O Key. With Key.
@@
 Average 92.4% 98.5% 64.4% 83.8%
-Experimental Results.
-We present the experimental results of utilizing Key Code Extraction on SeVC in Table IX and DiverseVul in Figure 6.
-We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type).
-For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code.
-The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul.
-The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks.
-This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models.
-Key code extraction offers two major benefits: (1) Enhanced Focus: By isolating the most relevant code segments, the model can focus on critical information, leading to more accurate detection results.
-(2) Noise Reduction: Eliminating irrelevant code improves the signal-to-noise ratio, enabling the model to better learn and detect vulnerability patterns.
-Moreover, this approach relies on automated annotation rather than manual feature extraction, making it broadly applicable and scalable.
-Our experimental results preliminarily validate the feasibility and utility of key code extraction through LLMs.
-Figure 6: Performance of Key Code on DiverseVul(RQ4).
-Summary of RQ4: By guiding LLMs to focus on key code, the performance of vulnerability detection can be significantly improved(e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
+TABLE XII: Impact of Key Code on Detection Performance on the PrimeVul Dataset (RQ4).
+Metric Without Key Code With Key Code
+Multi-Class CWE Detection
+Precision (Macro) 19.2 19.3
+Recall (Macro) 25.7 27.6
+F1 Score (Macro) 20.7 21.5
+Binary Vulnerability Detection
+Precision 29.5 31.7
+Recall 61.1 62.6
+F1 Score 39.8 42.1
+Figure 6: Impact of Key Code on Detection Performance on the DiverseVul Dataset (RQ4).
+Experimental Results. Our findings indicate that the effectiveness of key code extraction is highly dependent on the complexity of the dataset.
+On SeVC—the dataset featuring synthesized data and relatively simple vulnerability classification—key code extraction yields notably superior performance as demonstrated in Table XI.
+On the more challenging PrimeVul dataset, key code provides a clear advantage.
+As shown in Table XII, the “With Key Code” model consistently outperforms the baseline.
+The F1 score for binary detection improves from 39.8% to 42.1%, and the Macro-F1 for multi-class detection increases from 20.7% to 21.5%.
+This positive result is significant because PrimeVul is the benchmark where models struggled most (RQ1.4).
+The improvement stems from two benefits: (1) Enhanced Focus, which directs the model’s attention to the most relevant code, and (2) Noise Reduction, which improves the signal-to-noise ratio by filtering out irrelevant context.
+This demonstrates that for complex, noisy, real-world code, guiding the model’s focus is a highly effective strategy.
+On the DiverseVul dataset, the impact is inconclusive.
+In contrast, the results illustrated in Figure 6 show a mixed trend, with neither model establishing a definitive advantage.
+A plausible explanation is that the vulnerabilities in DiverseVul are more straightforward, allowing the baseline model to perform well without explicit guidance.
+In such cases, the overhead of the key code extraction process offers no significant additional benefit.
+Regarding explanation quality, our manual and LLM evaluation investigation that simply providing key code during fine-tuning did not yield a significant improvement in the quality of generated explanations.
+This suggests that a more sophisticated integration strategy is needed to translate focused attention into higher-quality explanatory text.
+Exploring how to better leverage key code to enhance explanation quality remains an important direction for future research.
+Summary of RQ4: Key code extraction enhances vulnerability detection performance on challenging real-world datasets by effectively guiding the model’s attention mechanisms.
 V Discussion
@@
 Implication 1: Broader Applications for Vulnerability-Related Tasks.
-Our study reveals that integrating vulnerability explanation with detection does not compromise the model’s performance.
-This finding suggests that we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs.
-By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights.
-For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime.
-This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
+Our finding that integrating vulnerability explanation with detection preserves model performance suggests expanding LLM fine-tuning beyond detection.
+Incorporating tasks like impact assessment, severity prioritization, or automated patch generation could enhance contextual understanding and actionability.
+This holistic approach may yield tools that simultaneously identify and mitigate vulnerabilities.
 Implication 2: Effectiveness of LLMVulExp Using Different LLMs.
 In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments.
-To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, to conduct the experiments following the setting of RQ1.3.
-Figure 7 shows the results of multi-type vulnerability detection using LLMs of different sizes.
-We find that LLMVulExp can achieve effective performance with these additional LLMs.
-Notably, the performance of CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct.
-Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
-Figure 7: Performance of different base LLMs.
+A key question is how our framework’s effectiveness varies with different base LLMs.
+We investigate this from both smaller and larger directions.
+First, we utilize two additional widely-used smaller LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, on the easier DiverseVul dataset.
+As shown in Figure 7, LLMVulExp achieves effective and comparable performance with these models, demonstrating that our proposed fine-tuning and explanation generation strategies are not limited to a single model architecture and can be successfully applied to smaller LLMs.
+Second, we seek to understand if simply employing a larger, more advanced model could further boost performance, particularly on a more challenging, realistic dataset.
+To this end, we conduct experiments on the PrimeVul dataset, comparing the performance of CodeLlama-13B-Instruct against a significantly larger and more powerful model, DeepSeek-Coder-33B-Instruct.
+The results, presented in Figure 8, lead to a crucial insight: the more advanced model did not yield a discernible accuracy gain.
+This finding consistently suggests that for fine-tuning tasks on current, small-scale vulnerability datasets, simply increasing the base model’s size or using a more recent architecture may not be the silver bullet.
+We hypothesize that the primary bottleneck for improving performance is now the scale and diversity of the training data, rather than the inherent capacity of the base model.
+This is a very valuable insight for the field, suggesting that future efforts should focus on data-centric approaches to unlock the full potential of large models for vulnerability detection.
+Figure 7: Different base LLMs on DiverseVul.
+Figure 8: Different base LLMs on PrimeVul.
 Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks.
-The effectiveness of fine-tuning heavily depends on the quality of the annotations used for training.
-Our study highlights the critical role that high-quality annotations play in the fine-tuning process.
-It is essential to develop robust annotation frameworks that ensure consistency and accuracy.
-Furthermore, the annotation process itself could be enhanced through the use of semi-automated tools that provide initial labels, which are then reviewed and refined by human experts.
-This hybrid approach could balance the need for detailed annotations and the practical constraints of manual labeling.
+To demonstrate the high sensitivity of fine-tuning effectiveness to annotation quality, we conduct experiments using different annotation LLMs.
+Specifically, CodeLlama-7B fine-tuned with DeepSeek-V3-generated explanations achieved an average Macro-F1 of 20.5% on PrimeVul, while the same model fine-tuned with CodeLlama-13B-generated explanations attained only 16.4%.
+This significant performance gap substantiates the critical role of annotation quality.
+This dependency underscores the necessity of developing robust annotation frameworks that ensure consistency and accuracy.
+A promising approach involves semi-automated tools where advanced LLMs provide initial annotations subsequently refined by human experts.
+Such a hybrid methodology could mitigate quality variations while balancing the practical constraints of manual curation.
 VI Threats to Validity
 Construct Validity.
-We evaluate explanatory capability using three effective metrics, but these metrics may not cover all aspects of explanations.
-To mitigate bias, we use both manual and automated reviews with LLMs.
-Two authors independently annotate the explanations, resolving discrepancies until consensus is reached, achieving a Cohen’s Kappa [46] value of 0.76, indicating substantial agreement.
+We evaluated explanatory capability using three metrics (Accuracy, Clarity, and Actionability).
+While informative, these metrics may not be exhaustive, and further refinement is desirable.
+To mitigate bias, two authors independently annotated explanations, achieving substantial agreement.
+Furthermore, our use of GPT-3.5 and DeepSeek-V3 for certain automated annotation tasks introduces a replication challenge: its specific version may deprecate or evolve, potentially impacting result reproducibility.
+Future work addressing this should meticulously document model versions and prompts or consider open-source alternatives.
+This overall rigorous approach supports our evaluation’s reliability, though we acknowledge this specific consideration for LLM-assisted annotation components.
 Internal Validity.
-Due to the high computational cost of fine-tuning and inference evaluation, we couldn’t conduct experiments under different data splits and randomization states.
-Although we limit the number of hyperparameter trials for LoRa configuration and didn’t fine-tune larger models like the 34B version due to GPU constraints, these limitations don’t undermine our objective: exploring the potential of LLMs for vulnerability detection and explanation.
+Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we are unable to conduct experiments under different data splits and randomization states.
+The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results.
+However, we conduct multiple trials to validate the stability of our conclusions for each experimental setup.
+Although the number of hyperparameter trials for LoRa configuration and the LLM generation method are relatively limited, and due to GPU constraints, we do not attempt to fine-tune larger models, we ensure that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
 External Validity.
-We choose the advanced open-source code LLM CodeLlama and evaluate different versions as well as the recently released Llama3 on two distinct datasets.
-These datasets, while containing a significant amount of vulnerable code, include only a few CWE types and mostly generic vulnerabilities, which do not fully reflect the diversity and complexity of vulnerabilities found in real-world development environments.
-Although these datasets contain substantial vulnerable code, they differ from real development environments, posing a risk of limited generalization.
+We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research.
+We test different versions, including 7B and 13B, and evaluate the performance of newer LLM Llama3 and DeepSeek-Coder on vulnerability detection and explanation tasks.
+However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments are limited to a few models.
+Additionally, we conduct experiments using three differently configured open-source project-based vulnerability datasets.
+Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments.
+Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments.
+Nonetheless, our careful selection and thorough evaluation process aim to ensure that our findings remain relevant and insightful within the scope of our study.
 VII Conclusion
@@
 Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.
+Acknowledgment
+This work was supported by the National Key R&D Program of China (No. 2024YFB4506400).